### v2

# Towards Architecture-Agnostic Untrained Network Priors for Image Reconstruction with Frequency Regularization

Abstract
Untrained networks inspired by deep image prior have shown promising capabilities in recovering a high-quality image from noisy or partial measurements, without requiring training data. Their success has been widely attributed to the spectral bias acting as an implicit regularization induced by suitable network architectures. However, applications of such network-based priors often entail superfluous architectural decisions, overfitting risks, and slow optimization, all of which hinder their practicality. In this work, we propose efficient, architecture-agnostic methods for a more direct frequency control over the network priors: 1) constraining the bandwidth of the white-noise input, 2) controlling the bandwidth of the interpolation-based upsamplers, and 3) regularizing the Lipschitz constants of the layers. We show that even with just one extra line of code, the overfitting issues in underperforming architectures can be alleviated such that their performance gaps with the high-performing counterparts can be largely closed despite their distinct configurations, mitigating the need for architecture tuning. This then makes it possible to employ a more compact model to achieve similar or superior performance to larger models with greater efficiency. Our regularized network priors compare favorably with current supervised and self-supervised methods on MRI reconstruction and image inpainting tasks, serving as a stronger zero-shot baseline reconstructor. Our code will be made publicly available.
1 Introduction
Magnetic resonance imaging (MRI) is a mainstream imaging tool for medical diagnosis. Reconstructing MR images from raw measurements refers to the transformation from Fourier spectrum of the object in $k$-space to image space. Since acquiring full $k$-space measurements is time-consuming, under-sampled $k$-space data are often collected to reduce scan times. Accelerated MRI is thus known as an ill-posed inverse problem that conventionally requires handcrafted priors [31, 27] to mitigate the resulting aliasing artifacts in the output images. While supervised learning methods based on convolutional neural networks (CNNs) have demonstrated better reconstruction quality with fewer measurements, their training relies on paired under-sampled and fully-sampled measurements, which are expensive to acquire and could raise issues on robustness and generalization when the acquisition protocol or anatomy changes [23, 24].
Figure 1: Example results from underperforming architectures with 4×4$\times$4 × under-sampling. Turning the left to the right simply by low-pass filtering the white-noise input via a Gaussian blur kernel, which can be implemented as few as one or two lines of code. SSIM (normal-$\uparrow$) values are reported.
Instead of requiring large-scale datasets for capturing prior statistics, the untrained network inspired by deep image prior (DIP) [45] requires only the corrupted or partial measurements and regularizes the reconstruction solely through the inductive biases of its architecture, making it a representative zero-shot self-supervised reconstruction method for a wide range of natural and medical applications [50, 28, 37]. Concretely, DIP parameterizes the unknown desired image via a neural network and optimizes the network parameters such that the output image transformed by the degradation model matches the acquired measurements. Such parameterization is observed to exhibit high resistance to noise and corruption, acting as a form of implicit regularization. Studies have attributed this property to CNN’s inherent spectral bias – the tendency to fit the low-frequency signals before the high-frequency signals (e.g., noise) [41, 5], where the choice of network architecture is often shown to be critically relevant [5, 30, 1].
Despite its great promise, favorable results obtained from the untrained network prior are contingent upon two critical factors: an optimal architecture for a specific task and an early-stopping strategy to prevent overfitting to the noisy, partial measurements. Furthermore, while being agnostic to domain shifts by optimizing on a per-image basis [48], such reconstruction process is inherently slow, hindering its practicality. While these issues are often intertwined — inappropriate and over-parameterized architectures usually exacerbate overfitting and increase runtime, most prior works tackle these challenges separately. For architectural design, existing methods rely on either handcrafting or utilizing neural architecture search techniques [14, 9, 1, 6, 26, 16]. However, the lack of a consensus on the desired architectural prior often leads to a laborious search. Another line of work is dedicated to preventing overfitting through early-stopping [48, 46], subspace optimization [2] or pretraining-then-fine-tuning [3, 35], which also helps shorten the runtime. Similar in spirit to traditional ways of combating overfitting, these methods aim to use fewer trainable parameters with little performance degradation (and may involve costly pretraining) or a held-out subset of measurements for self-validation.
In this work, we explore the possibility of modulating the frequency bias and hence the regularization effects of network priors in an architecture-agnostic manner, aiming to enhance a given architecture irrespective of its specific configurations. This is conceivable, in light of the recent body of theoretical and empirical evidence indicating that there are only a few key componenets within the architecture that are driving forces behind the spectral bias in DIP [30, 5, 41, 15]. Inspired by these findings, we develop simple yet effective methods from a frequency perspective to efficiently regularize the network priors, alleviating overfitting by curbing the over-fast convergence of high frequencies, all with minimal architectural modifications and computational costs. Specifically, we propose to (i) constrain the effective bandwidth of the white-noise input via simple spatial blurring or selected Fourier features, (ii) adjust the bandwidths of the interpolation-based upsamplers to control their attenuation (smoothing) extents, and (iii) regularize the Lipschitz constants of the layers to enforce function smoothness. We empirically find that our regularized network priors not only exhibit less vulnerability to overfitting but also tend to achieve better extrapolation and thus higher peak performance in inpainting-like tasks. Demonstrated in the context of MRI reconstruction, our methods significantly improve models across various architectural configurations, largely reducing the need for extensive architectural tuning (Fig. 1). The efficacy of our methods is also showcased in image inpainting and denoising tasks, where the control over frequency is adjusted accordingly. By minimizing architectural influences, our approach additionally offers a unique advantage in efficiency: a smaller, previously underperforming network, can now achieve on-par performance or even surpass a larger, heavily parameterized high-performing network. Our contributions are three-fold:
  • We propose efficient methods that directly modulate the frequency bias in the untrained network priors in an architecture-agnostic manner, effectively addressing challenges related to architecture, overfitting, and runtime in a unified approach.
  • The enhanced untrained networks match the performance of leading self-supervised methods and surpass the supervised method especially on out-of-domain data, with up to 86×86$\times$86 × reduction in runtime ($\sim 5$ mins/slice) for MRI reconstruction, serving as a stronger zero-shot reconstruction baseline than the original DIP scheme.
  • Our empirical findings and results on medical and natural image reconstruction tasks shed light on the spectral behaviors of network-based priors—CNNs operating in a single-instance generative setting.
2 Related Work
Function Smoothness and Spectral Bias. Function smoothness, also referred to as function frequency, quantifies how much the output of a function varies with changes in its input [12]. Spectral bias [38, 47] is an implicit bias that favors learning functions changing at a slow rate (low-frequency), e.g., functions with a small Lipschitz constant. In visual domains, this is evident in the network’s output lacking subtle details. Many regularization techniques shown to aid generalization encourage smoothness implicitly, such as early stopping, $\ell_{2}$ regularizer [39]. Smoothness has been widely used as a model complexity measure in place of model size to account for the well-known "double-descent" phenomenon associated with generalization [34]. To explicitly promote smoothness, it is natural to penalize the norm of the input-output Jacobian [36, 17]. However, due to the high dimensionality of the output such as in accelerated MRI, computation of the Jacobian matrix during training is often intractable. Another efficient and prevalent solution is to constrain the network to be c-Lipschitz with a pre-defined Lipschitz constant c [33, 13]. We develop a suitable form of Lipschitz regularization for untrained networks by instead penalizing learned Lipschitz constants, with a novel aim of achieving architecture-agnostic untrained image reconstruction.
Input Frequency and Generalization. Input has played an important role in helping neural networks represent signals of various frequencies. As in implicit neural representations (INRs) [42] and neural radiance fields (NeRFs) [32] where coordinates are mapped to RGB values, naively training with raw coordinates as inputs results in over-smoothing; encoding the input coordinates with sinusoidal functions of higher frequencies enables the network to represent higher frequencies [32, 43]. Rahaman et al. [38] also show theoretically and empirically that fitting becomes easier for the network when the input itself contains high-frequency components. However, it has recently been reported that the high-frequency positional input encodings lead to the failure of NeRFs in few-shot settings due to over-fitting [49]. Here, we show that similar issue also applies to untrained network priors and can be efficiently addressed by our methods.
Avoid Overfitting in Untrained Networks. A line of effort has been exclusively devoted to preventing overfitting to noisy images or measurements. Wang et al. [46] propose to track the running variance of the output for an early-stopping criterion, but it is found to be unstable in medical image reconstruction [2]. Yaman et al. [48] propose to split the available measurements into a training and a validation subset and use the latter for self-validation for automated early stopping. Transfer-learning-based untrained networks perform pre-training on synthetic data followed by fine-tuning [3, 35] or subspace optimization [2]. These methods aim to use fewer trainable parameters to avoid overfitting with little performance degradation. In contrast, our methods alleviate overfitting from a frequency perspective and enable significantly better performance while maintaining the same or less model-wise and computation-wise complexity.
Figure 2: Overview of the proposed regularized network priors for MRI reconstruction.
3 Method
3.1 Preliminaries
Accelerated MRI The goal of accelerated MRI reconstruction is to recover a desired image $\mathbf{x\in\mathbb{C}}^{n}$ ($n=n_{h}\times n_{w}$) from a set of under-sampled k-space measurements. We focus on a multi-coil scheme in which the forward model is defined as:
$\mathbf{y}_{i}=\mathbf{A}_{i}\mathbf{x}+\epsilon\text{ with }\mathbf{A}_{i}=\mathbf{MFS}_{i},\quad i=1,\dots,c,$ (1)
where $\mathbf{y}_{i}\in\mathbb{C}^{m}$ denotes the $k$-space measurements from coil $i$, $c$ denotes the number of coils, $\mathbf{S}_{i}\in\mathbb{C}^{n}$ denotes the coil sensitivity map (CSM) that is applied to the image $\mathbf{x}$ through element-wise multiplications, $\mathbf{F}\in\mathbb{C}^{n\times n}$ denotes the 2D discrete Fourier transform, $\mathbf{M}\in\mathbb{C}^{m\times n}$ denotes the under-sampling mask, and $\epsilon\in\mathbb{C}^{m}$ denotes the measurement noise.
Untrained MRI Reconstruction is often framed as an inpainting problem where the network recovers the unacquired $k$-space measurements (masked) based on the acquired $k$-space data (observed). The image $\mathbf{x}$ is parameterized via a neural network $\mathbf{G_{\theta}(z)}$ with a fixed noise input vector $\mathbf{z}$ drawn from a uniform distribution $z\sim\mathcal{U}(0,1)$. With the aforementioned MRI forward model, the untrained network solves the following optimization problem:
$\theta^{\ast}=\mathop{\arg\min}\limits_{\theta}\mathcal{L}(\mathbf{y};\mathbf{AG_{\theta}(z)}),\quad\mathbf{x}^{\ast}=\mathbf{G_{\theta^{\ast}}(z)}.$ (2)
Such parameterization allows novel image priors to be designed dependent on the network architecture and the associated parameters. Nevertheless, many studies augment the untrained networks with traditional image regularizers [29], e.g., Total variation (TV), though it can only partially alleviate over-fitting [35, 2]. In our experimental sections, we show that TV is not as effective as ours in improving under-performing architectures.
3.2 Architecture-agnostic Frequency Regularization
To modulate the regularization effects of the network prior, we identify three sources of spectral bias within the network and introduce corresponding regulation methods, as depicted in Fig. 2. Each method requires only minimal hyperparameters to achieve architecture-agnostic control over the network priors.
Figure 3: (a) Visualization of the 1D white noise, the low-passed noise via Gaussian blur, and Fourier features of various frequencies in the frequency domain. (b) Gaussian blurred noise input and Fourier features with lower $f_{c}$ ($L=4$ or $8$) are both effective in alleviating overfitting and enhancing the peak performance.
Bandwidth-constrained Input. An important aspect of untrained networks that can be easily overlooked is their inputs. Conventionally, the inputs are randomly sampled from either a uniform or Gaussian distribution and are then mapped to an image. From a frequency perspective, such white-noise input comprises all frequencies with uniform intensities [11]. With this view in mind, we draw an analogy between untrained networks and INRs that map Fourier features to RGB values. Fourier features are sinusoid functions of the input coordinates $\mathrm{\mathbf{p}}$, i.e., $[\sin(\mathrm{2^{0}\pi\mathbf{p}}),\cos(\mathrm{2^{0}\pi\mathbf{p}}),...,\sin(2^{L-1}\pi\mathbf{\mathrm{\mathbf{p}}}),\cos(2^{L-1}\pi\mathbf{\mathrm{\mathbf{p}}})]$, where a larger $L$ assists the network in representing higher-frequency functions [43].
Figure 4: Narrowing input’s bandwidth enhances performance as a "free lunch". The output becomes smoother as $\sigma$ increases, up to a certain point.
White noise with variance $\sigma^{2}$ exhibits an autocorrelation that is a scaled Dirac $\delta$-function $\sigma^{2}\delta(t)$, whose Fourier transform $\mathcal{F}$ has a constant magnitude $\sigma^{2}$ spanning all frequencies $\mu$, i.e., $\mathcal{F}\{\sigma^{2}\delta(t)\}(\mu)=\sigma^{2}$, $\mu\in\mathbb{R}$.
In this sense, an untrained network can be viewed as mapping a broad spectrum of Fourier features to a target image (Fig. 3 (a)). We hypothesize that this enhances the network’s representation ability but likely incurs over-fitting due to the faster convergence of high-frequency components. To validate this, we applied a Gaussian blur filter $\mathcal{G}_{s,\sigma}$ on the white-noise input $z$ to remove a certain amount of high frequencies before passing it to the network, defined as $z*\mathcal{G}_{s,\sigma}$, where *** denotes convolution. The filter size $s$ and sigma value $\sigma$ that controls the filter’s bandwidth are hyperparameters. As exemplified in Fig. 4, simply adjusting $\sigma$ already brings significant gains without architectural changes. Similarly, substituting the noise input with Fourier features, with a carefully selected maximum frequency $f_{c}\propto L$ to limit the input’s effective bandwidth has also proven effective in Fig. 3 (b) and our ablation study (Tab. 23), further validating our hypothesis. Fourier features introduce frequency-diverse input akin to white noise but in a controlled manner, enabling frequency regularization.
Drawing from these empirical findings, we propose to regularize the input’s frequency range, achievable via simple spatial blurring or selected Fourier features, to efficiently mitigate overfitting and enhance network performance. While Gaussian blur and Fourier Features both offer bandwidth control, they yield different frequency representations (e.g., Fourier features are more discrete than Gaussian blurred noise), which may affect performance and susceptibility to parameter tuning differently. We present and compare both results in the experiments.
We note that inputs have been deemed crucial in guiding the generation of the final image not only in INRs but also in generative networks, e.g., transformed Fourier features have been used as an alternative to random input in StyleGAN3 [21], for a different goal of improving the architecture’s equivariance. In contrast, we use a low-passed version of Fourier features or noise input to prevent over-fast convergence on high frequencies and enable better extrapolation. Similar in spirit, recent work by [49] also shows that masking the high-frequency Fourier features helps NeRFs generalize in few-shot settings.
Table 1: Influences of upsamplers on reconstruction. From the left to the right, the attenuation extent of the upsampling method increases. Construction details of the upsampler $\mathcal{L}_{-90}$ follows [30]. Frequency responses of the interpolation filters are shown in the figure below. Evaluated on the 4×4$\times$4 × fastMRI multi-coil brain datasets.
Methods w/o. Upsampling. Nearest Bilinear $\mathcal{L}_{-90}$ of Params. (Millions) ConvDecoder [9] 28.69 $\pm$ 1.6 31.78 $\pm$ 1.2 32.31 $\pm$ 1.3 32.48 $\pm$ 1.2 4.1 M Deep Decoder [14] 24.55 $\pm$ 1.1 27.10 $\pm$ 0.9 31.36 $\pm$ 1.4 32.68 $\pm$ 1.1 0.47 M
Smoothness-controlled Upsampling. We found that constraining the input’s bandwidth significantly enhances shallower architectures, yet the improvement diminishes as the network depth increases. This could partly be attributable to the fact that subsequent layers along with nonlinearity can always generate new arbitrarily high frequencies, as theoretically and empirically proved in [38, 21].
We notice that the interpolation-based upsampling methods within the network, such as nearest neighbor and bilinear, are essentially implicit low-pass filters smoothing out the alias frequencies caused by the increased sampling rate. Prior works [30, 15] on image denoising have shown that these non-trainable upsampling methods are driving forces behind the spectral bias of the network priors, delaying the convergence of higher frequencies. Different upsamplers bias the network towards different spectral properties [30]. We confirm that upsampling also substantially influences image reconstruction and further reveal that an appropriate upsampling method can bridge the performance gap between two distinct architectures (Tab. 1).
Motivated by these results, we introduce a controllable upsampler to modulate the network’s spectral bias, especially the deeper architectures. We achieve this by first interleaving the input feature maps with zeros, and then convolving them with a customized low-pass filter with adjustable bandwidth (Fig. 2). For filter design, we adopt the Kaiser-Bessel window as it offers explicit control over the tradeoffs between passband ripple and stopband attenuation. The Kaiser window is defined as
$w(n)=I_{0}(\beta\sqrt{1-(2n/M)^{2}})/I_{0}(\beta),-M/2\leq n\leq M/2,$ (3)
where $M$ is the desired spatial extent of the window, $\beta\geq 0$ is the shape parameter—the higher it is, the greater the stopband attenuation is (and generally the smoother the image is), and $I_{0}$ is the zeroth-order modified Bessel function of the first kind [20]. This plug-and-play upsampler can be inserted in different layers with different $M$ and $\beta$ hyperparameters, offering flexible and precise control.
Lipschitz-regularized Layers. Spectral bias towards low frequencies favors functions that do not change at a high rate, i.e., functions with small Lipschitz constants. A function $f:\mathcal{X}\rightarrow\mathcal{Y}$ is said to be Lipschitz continuous if there is a constant $k>0$ such that
$\|f(\mathrm{x_{1}})-f(\mathrm{x_{2}})\|_{\text{p}}\leq k\|\mathrm{x_{1}}-\mathrm{x_{2}}\|_{\text{p}}\quad\forall\mathrm{x_{1}},\mathrm{x_{2}}\in\mathcal{X},$ (4)
where $k$ is the Lipschitz constant that bounds how fast $f$ can change globally w.r.t. input perturbations.
Instead of upper bounding the Lipschitz constants of the network layers to pre-defined and manually chosen values as in [41], we make the per-layer Lipschitz bounds learnable and regularize their magnitudes during optimization.
The Lipschitz constant of a convolutional layer is bounded by the operator norm of its weight matrix [13]. To bound a convolutional layer to a specific Lipschitz constant $k$, the layer with $m$ input channels, $c$ output channels and kernels of size $w\times h$ is first reshaped to a 2-D matrix $W^{\ell}\in\mathbb{R}^{n\times cwh}$, and then normalized as:
$\tilde{W_{\ell}}=\frac{W_{\ell}}{\text{max}(1,\frac{\|W_{\ell}\|_{\text{p}}}{\mathtt{SoftPlus}(k_{\ell})})},$ (5)
where $k_{\ell}$ is a learnable Lipschitz constant for each layer, $\|\cdot\|_{\text{p}}$ is chosen as the $\ell_{\infty}$ norm and $\mathtt{SoftPlus}(k_{\ell})=\text{ln}(1+\text{exp}(k_{\ell}))$ ensures the learned Lipschitz bounds are non-negative. Such formulation only normalizes $W_{\ell}$ if its matrix norm is larger than the learned Lipschitz constraint during training. Integrating the ultimate Lipschitz regularization with learned Lipschitz constants into Eq. 2, our regularized training objective is:
$\min_{\Theta,K}\mathcal{L}(\mathbf{y};\mathbf{AG_{\Theta}(z)})+\lambda\sum_{l=1}^{L}\mathtt{SoftPlus}(\mathbf{k}_{\ell})^{2}$ (6)
where $\mathbf{K}$ is a collection of per-layer learnable Lipschitz constant $k_{\ell}$ jointly optimized with the network parameters, and $\lambda$ controls the granularity of smoothness.
Figure 5: Our approach significantly minimizes the performance gaps among architectures with various depths $\{2,5,8\}$ and widths $\{64,256\}$.
4 Experiments
4.1 Setup and datasets
We first (1) validate the proposed methods’ effectiveness in enhancing the performance of untrained networks across various architectural configurations, especially those originally underperforming. We then (2) benchmark the enhanced versions of those compact architectures against established supervised and self-supervised methods on both in-domain and out-of-domain datasets in terms of accuracy and efficiency. We also (3) compare our methods with self-validation-based early stopping [48] on overcoming overfitting, and show that it is complementary to our approach by further shortening the reconstruction time. Finally, we demonstrate the utility of our methods in (4) general image inpainting and denoising tasks and perform spectral bias analysis on all evaluated tasks.
The MRI experiments were performed on two publicly available datasets including the multi-coil knee and brain MRI images from fastMRI database [25], and multi-coil knee MRI images from Stanford 3D FSE knee dataset [19]. The fully-sampled $k$-space data was retrospectively masked by selecting 25 central $k$-space lines along with a uniform undersampling at outer $k$-space, achieving the standard 4×4$\times$4 × acceleration. For training the supervised baseline, the knee training set consists of 367 PD and PDFS slices and the brain training set consists of 651 slices with a mixture of T1, T2 weighted images. 50 knee slices and 50 brain slices were sampled from the respective multi-coil validation datasets for evaluation.
4.2 Implementation Details
Without loss of generality, the base architectures considered in our work are N-level encoder-decoder architectures with full skip connections. The architectures are isotropic with the same width and kernel size throughout the layers. All evaluated architectures are trained for 3k iterations using mean absolute error and Adam optimizer [22] with a learning rate of 0.008. Unless otherwise specified, the results at the last iteration are reported. The input is drawn from a uniform distribution $z\sim\mathcal{U}(0,1)$. The filter size of the Gaussian blur was set to 5 and the sigma value was randomly sampled from $[0.5,2.0]$ for every slice. $M$ and $\beta$ for the Kaiser-based upsamplers are chosen to be $\{15\times N-1,5\}$ and $\{5\times N\}$ for knee data ($N$ denotes the nth-level), and $\{5\times N\}$ and $\{5\times N\}$ for brain data, respectively. $\lambda$ is set to $1$ for Lipschitz regularizer.
Table 2: Effectiveness of the methods in bridging performance gaps among different architectures, evaluated on fastMRI brain datasets. Bandlimited inputs achieved by Fourier features ($L=4$ or $8$) or Gaussian blur along with Lipschitz regulariation improve all architectures, especially the shallower. The proposed Kaiser-based upsampling dramatically improves the deeper architectures. All architectures end up with similarly high performance. The best and the second-best are highlighted.
Regularizers $\mathbf{A_{2}\mathunderscore_{256}}$ $\mathbf{A_{2}\mathunderscore_{64}}$ $\mathbf{A_{5}\mathunderscore_{256}}$ $\mathbf{A_{5}\mathunderscore_{64}}$ $\mathbf{A_{8}\mathunderscore_{256}}$ $\mathbf{A_{8}\mathunderscore_{64}}$ $\mathbf{A_{2}\mathunderscore_{256}}$ $\mathbf{A_{2}\mathunderscore_{64}}$ $\mathbf{A_{5}\mathunderscore_{256}}$ $\mathbf{A_{5}\mathunderscore_{64}}$ $\mathbf{A_{8}\mathunderscore_{256}}$ $\mathbf{A_{8}\mathunderscore_{64}}$ PSNR $\uparrow$ SSIM $\uparrow$ w/o. Reg. (Plain) 29.08 29.41 31.15 31.42 31.27 31.68 0.729 0.761 0.782 0.801 0.784 0.807 TV 29.22 29.61 31.26 31.37 31.32 31.64 0.735 0.764 0.785 0.802 0.787 0.807 Lipschitz Reg. 30.92 29.73 31.47 32.11 31.50 32.03 0.795 0.766 0.792 0.812 0.800 0.820 Fourier features ($L=16$) 30.57 30.49 31.57 31.77 31.77 32.09 0.786 0.788 0.794 0.813 0.799 0.819 Fourier features ($L=8$) 31.42 31.98 31.82 32.42 31.60 32.45 0.804 0.833 0.799 0.831 0.795 0.834 Fourier features ($L=4$) 31.92 32.59 31.87 32.80 31.71 32.86 0.840 0.863 0.799 0.848 0.793 0.844 Gaussian blurred 33.34 32.67 32.14 32.66 32.03 32.92 0.870 0.866 0.811 0.849 0.825 0.849 Gauss. + Lips. 32.90 33.12 32.08 32.83 31.70 33.14 0.855 0.870 0.815 0.851 0.805 0.849 Gauss. + Lips. + Kaiser Up. 32.50 33.10 33.00 33.21 33.09 33.85 0.836 0.874 0.857 0.876 0.858 0.885
Table 3: Evaluated on fastMRI knee datasets.
Regularizers $\mathbf{A_{2}\mathunderscore_{256}}$ $\mathbf{A_{2}\mathunderscore_{64}}$ $\mathbf{A_{5}\mathunderscore_{256}}$ $\mathbf{A_{5}\mathunderscore_{64}}$ $\mathbf{A_{8}\mathunderscore_{256}}$ $\mathbf{A_{8}\mathunderscore_{64}}$ $\mathbf{A_{2}\mathunderscore_{256}}$ $\mathbf{A_{2}\mathunderscore_{64}}$ $\mathbf{A_{5}\mathunderscore_{256}}$ $\mathbf{A_{5}\mathunderscore_{64}}$ $\mathbf{A_{8}\mathunderscore_{256}}$ $\mathbf{A_{8}\mathunderscore_{64}}$ PSNR $\uparrow$ SSIM $\uparrow$ w/o. Reg. (Plain) 27.18 27.62 29.16 29.23 28.98 29.35 0.541 0.575 0.628 0.640 0.625 0.644 TV 28.25 27.85 29.33 29.57 29.54 30.01 0.588 0.592 0.635 0.651 0.645 0.687 Lipschitz Reg. 28.41 29.21 29.17 29.79 29.43 30.14 0.601 0.600 0.629 0.651 0.636 0.666 Fourier features ($L=16$) 28.42 28.97 29.58 30.26 29.76 30.38 0.587 0.622 0.653 0.671 0.661 0.681 Fourier features ($L=8$) 28.61 29.98 29.86 30.72 29.66 30.89 0.604 0.670 0.669 0.693 0.662 0.703 Fourier features ($L=4$) 32.02 32.07 29.40 31.13 29.55 31.17 0.775 0.781 0.665 0.718 0.668 0.717 Gaussian blurred 30.87 30.89 30.02 31.24 29.31 30.89 0.739 0.768 0.694 0.748 0.698 0.727 Gaussian blurred + Lips. 31.61 31.93 29.40 31.67 29.82 31.58 0.750 0.776 0.702 0.727 0.697 0.732 Gauss. + Lips. + Kaiser Up. 31.92 31.61 31.78 31.60 31.09 31.73 0.777 0.776 0.778 0.776 0.750 0.768
4.3 Effectiveness in reducing architectural sensitivity
Fig. 5 gives a quantitative overview of the substantial improvement using our approach in architectures with diverse configurations on both knee and brain datasets. The different results of the original architectures also confirm the influences of architectural choices on performance. Notably, before applying our methods, the deeper and narrower architectures tend to perform better than their counterparts (more in appendix). This trend aligns with previous works [9, 3, 10, 45] where these architectures tend to be favored in inpainting-like tasks. Here we identify particularly their counterparts (i.e., $\mathbf{A_{x}\mathunderscore_{256}}$ ) as "underperforming" architectures. As will be shown in our spectral bias analysis in Sec. 4.5 and appendix, these underperforming architectures learn high frequencies more quickly (though this may be desired for other tasks [30]) and are more susceptible to overfitting, incurring severe artifacts in the output (Figs. 6 and 9). When applied with our methods, as detailed in Tab. 2 and Tab. 3, a large boost in performance is observed in all architectures, especially in $\mathbf{A_{2}\mathunderscore_{256}}$.
We observe that using low-passed inputs via either selected Fourier features or blurring brings the most benefits to the shallower architectures. Better results are achieved when combined with Lipschitz regularization on the layers. On the other hand, deeper architectures benefit more from the Kaiser-based upsampler, which can be seen as performing low-pass filtering on the input feature maps within the network, beyond the initial input layer. We further note that the hyperparameters required for upsampling differ between knee and brain data (Sec. 4.2), with the knee data requiring greater attenuation. This is also consistent with previous findings that they require different numbers of channels for the architecture [9]. Our methods greatly alleviate the need for such architectural tuning by instead allowing for the adjustment of a few key hyperparameters.
Figure 6: Our methods enable the underperforming architectures (e.g., $\mathbf{A_{8}\mathunderscore_{256}}$, $\mathbf{A_{2}\mathunderscore_{256}}$) to perform similarly to the well-performing architectures (e.g., $\mathbf{A_{8}\mathunderscore_{64}}$).
Table 4: Quantitative results on fastMRI datasets. Runtime: mean (std) per slice.
Datasets Supervised UNet CS-$\ell_{1}$ [18] ZS-SSL [48] DIP [45] Deep Decoder [14] ConvDecoder [9] $\mathbf{A_{2}\mathunderscore_{64}}$ (vanilla) $\mathbf{A_{8}\mathunderscore_{64}}$ (vanilla) $\mathbf{A_{2}\mathunderscore_{64}}$ (Ours) $\mathbf{A_{8}\mathunderscore_{64}}$ (Ours) Brain PSNR $\uparrow$ 33.35 29.91 34.39 31.15 26.97 31.81 29.42 31.68 33.10 33.85 SSIM $\uparrow$ 0.889 0.773 0.878 0.782 0.747 0.800 0.761 0.807 0.874 0.885 Knee PSNR $\uparrow$ 31.15 28.23 32.00 29.16 27.21 29.59 27.62 29.35 32.07 31.73 SSIM $\uparrow$ 0.776 0.633 0.773 0.628 0.687 0.655 0.575 0.644 0.781 0.768 GFLOPS $\downarrow$ 99.24 5461.6 615.72 82.82 699.94 38.42 40.94 62.36 68.38 Runtime (mins) $\downarrow$ 0.002 (0.00003) 64.8 (20.18) 14.0 (0.61) 6.6 (0.63) 8.2 (0.35) 5.4 (0.47) 10.5 (0.62) 6.6 (0.58) 12.3 (0.65)
Figure 7: Qualitative evaluations. In-domain: fastMRI. Out-of-domain: Stanford FSE.
Table 5: Out-of-domain evaluation among supervised and untrained methods and comparisons with self-validation-based early stopping[48].
Method In-domain Out-domain Runtime (mean±$\pm$±std) PSNR SSIM PSNR SSIM Train Inference Trained U-Net 31.16 0.776 29.16 0.724 $\sim$1.5 days $0.1\pm 0.003$ sec Untrained CS-$\ell_{1}$ [18] 28.23 0.633 22.46 0.407 ZS-SSL [48] 32.00 0.773 31.74 0.805 $26.1\pm 3.5$ mins DIP [45] 29.16 0.628 28.89 0.664 $9.2\pm 0.3$ mins $\mathbf{A_{2}\mathunderscore_{64}}$ 27.62 0.575 26.03 0.550 $5.5\pm 0.1$ mins $\mathbf{A_{2}\mathunderscore_{64}}$ w. Self-Val 29.59 0.695 27.59 0.641 $0.2\pm 0.2$ mins $\mathbf{A_{2}\mathunderscore_{64}}$ (Ours) 32.07 0.781 31.43 0.790 $6.4\pm 0.4$ mins $\mathbf{A_{2}\mathunderscore_{64}}$ (Ours) w. Self-Val. 31.97 0.776 31.30 0.800 $0.3\pm 0.1$ mins
4.4 Benchmark results
We then compare our regularized network priors with several established MRI reconstruction methods, including a supervised baseline, a state-of-the-art self-supervised method (ZS-SSL) [48] and other untrained networks (ConvDecoder [9], Deep Decoder [14]). We adopt two lightweight architectures with distinct configurations as the base model. Visual comparisons are shown in Fig. 7.
Our methods enable the previously underperforming architectures to match the performance of ZS-SSL and the supervised UNet on in-domain fastMRI knee data (Tab. 4) and surpass the trained UNet on out-of-domain Stanford 3D FSE data (Tab. 5) as untrained methods are agnostic to domain shifts. The datasets for training the UNet contain 367 knee slices and 651 brain slices, respectively, covering various kinds of contrast weightings. Our enhanced networks also clearly outperform other untrained networks and the compressed-sensing (CS) based reconstruction. These results demonstrate the benefits of enhancing the untrained methods for generalizable reconstruction. ZS-SSL is an unrolling-based method where the network, i.e., a ResNet [44], is adopted as a denoiser. Comparisons with its UNet-variant, i.e.,ZS-SSL-UNet, are included in the appendix. Additionally, ZS-SSL uses a self-validation strategy for early stopping. We further compare it with our methods. Tab. 5 shows that self-validation is not as effective as ours. This is because, while self-validation mitigates overfitting by stopping near the peak PSNR, it does not fundamentally address the issues of underperforming architectures, i.e., the peak PSNR remains subpar. Our methods alleviate overfitting while generally enhancing the peak performance (e.g., Fig. 3), which we further analyze in the next section. Hence, self-validation-based early stopping is complementary to our goal of reducing the architectural sensitivity, and can be integrated into our approach to further shorten the reconstruction time as shown in Tab. 5. Overall, our enhanced network priors achieve comparable performance than ZS-SSL while being orders of magnitude faster.
4.5 Spectral bias analysis
Figure 8: Measurement of the spectral bias. Our methods allow high frequencies to be fitted at a later iteration, alleviating overfitting and generally beneficial to inpainting-like tasks (i.e., MRI) with higher peak performance. PSNR shown for $\mathbf{A_{2}\mathunderscore_{256}}$.
To examine the efficacy of the proposed methods on frequency control over the network, we measure the spectral bias using the metric— frequency-band correspondence (FBC) [41] that calculates the element-wise $|\mathcal{F}(x)|/|\mathcal{F}(y)|$ between the output $x$ and target image $y$, categorizes it into five frequency bands radially and computes the per-band averages. Higher values indicate higher correspondence. We trace the evolutions of FBC for $\mathbf{A_{8}\mathunderscore_{256}}$ and $\mathbf{A_{2}\mathunderscore_{256}}$ and the corresponding PSNR curves throughout the training iterations in three different tasks.
As shown in Fig. 8, the original architectures tend to readily fit all frequencies, especially $\mathbf{A_{2}\mathunderscore_{256}}$, corresponding to its worst performance among all compared architectures. Our methods delay the convergence of higher frequencies in both architectures across all three tasks as designed, which generally leads to higher peak performance for inpainting-like tasks. We hypothesize that a bias towards lower frequencies helps the model leverage spatial information for improved inter/extrapolation capabilities. Qualitative results in Fig. 9 validate our hypothesis. Given that MRI reconstruction resembles an inpainting task for $k$-space measurements, we expect similar improvement using our regularized network priors, as proved in our experiments. Note that although a stronger spectral bias may not always be optimal, the hyperparameters of our methods enable flexible adaptations for the specific task (more experiments in the appendix).
Figure 9: Inpainting and denoising ($\sigma=25$) experiments for demonstrating the improvement on underperforming architectures.
5 Conclusion
We introduce efficient, architecture-agnostic methods for frequency control over the network priors, offering a novel solution to simultaneously address the key challenges present in untrained image reconstruction. Our approach requires only minimal modifications to the original DIP scheme while achieving significant gains in accruacy and efficiency as evidenced in MRI reconstruction and natural image restoration tasks, making it a stronger zero-shot baseline reconstructor with the potential for seamless integration with other advancements in self-supervised reconstruction.
Appendix
A Natural image experiments
A.1 Inpainting
To the best of our knowledge, our method is the first attempt that addresses challenges related to architecture, overfitting and runtime simultaneously.
To compare with prior DIP methods, we 1) first employ the network configurations that work best in their respective settings [7, 41], i.e., the architectures used are different in different methods, as shown in Fig. 10 and Fig. 11, we then 2) use an underperforming architecture for all competing methods, as shown in Fig. 12.
Figure 10: Qualitative comparisons with previous DIP methods on inpainting.
Figure 11: Qualitative comparisons with previous DIP methods.
Figure 12: When unsuitable architectures are used, e.g., UNet w. deconvolutions, ResNet w/o. upsampling, SGLD [7] fails to perform restoration. This confirms the importance of architectural decisions in DIP, and that previous methods do not address the architecture-related challenges. Note that deconvolutions have been reported to be not suitable for DIPs [30, 15]. Similarly for ResNet, which does not have any upsampling layers [30].
A.2 Denoising
We first show the results when each method operates in its original setting (Fig. 13), and then evaluate them when an unsuitable architecture is used (Fig. 14).
Figure 13: Qualitative comparisons with previous DIP methods on denoising. Ours trades off the metrics for sharpness.
Figure 14: Denoising results. The base network used in each competing method is replaced by an underperforming architecture, i.e., $\mathbf{A_{2}\mathunderscore_{256}}$.
A.2.1 Transformer.
Figure 15: Qualitative denoising results of a transformer [4]. Our method substantially alleviates the overfitting and enhances the peak PSNR.
Besides CNN, we show here the result on Swin U-Net [4], which consists of only Swin Transformer blocks and skip connections, i.e., no upsampling is involved. As noted in a recent study [30], the unlearnt upsampling is the driving force behind the spectral bias of DIP, and such transformers are more difficult to perform denoising. This finding assumes the white noise as the input. Here, we show that constraining the bandwidth of the noise input enables long-lasting denoising even in such models.
B Influences of hyperparameters
As our methods contain several hyperparameters, we visualize their influences on the frequency control and hence the regularization effects over the output image in Fig. 16.
$s$ and $\sigma$ are associated with the Gaussian blur kernel applied on the noise input. The larger the $s$ and/or $\sigma$ is, the more high frequencies are removed from the noise input (i.e., the smoother the output is). $M$ and $\beta$ are for adjusting the attenuation extents of the Kaiser-based upsamplers. The higher the $M$ and/or $\beta$ is, the larger the attenuation of the high-frequency replica (i.e., the smoother the output is).
Figure 16: Visualizations of the frequency control with our methods in denoising experiments.
C Comparisons with early-stopping
Figure 17: Comparisons with self-validation-based early stopping. Although early stopping alleviates overfitting, it cannot fundamentally improve the underperforming architectures. Our methods can improve their peak performance while also mitigating overfitting.
To complement Table 5 in the main text, here we visualize in Fig. 17 that early stopping, even though prevents further performance decay, cannot fundamentally improve the underperforming architectures. In other words, early stopping could not cope with ill-designed architectures.
D Architectural influences
To better inform our method design, we 1) investigate the architectural influences in the context of MRI reconstruction, and also 2) validate our findings on image inpainting and denoising. Our results confirm that the reconstruction outcome is sensitive to basic architectural properties.
D.1 Crucial Architectural Elements
We first pinpointed the core architecture elements that have a critical impact on the performance.
Experimental setup i. Since a decoder is the minimum requirement for reconstruction, we experimented with two types of 7-layered decoder-only architectures, i.e., ConvDecoder [9] and Deep Decoder [14]. Experiments were performed on the 4× under-sampled multi-coil knee MRI from fastMRI database [25].
Figure 18: Influences of architectural elements. Results averaged across three different widths.
Upsampling (interpolation filter). Fig. 18(a) suggests an interesting result: removing the unlearnt upsampling, e.g., bilinear, leads to either failure or unstable results (see gray curves). Unlike transposed convolution, the unlearnt upsampler is essentially a fixed low-pass interpolation filter that attenuates the introduced high-frequency replica and also the signal. Frequency response of bilinear interpolation filter decays more rapidly than that of nearest neighbor as the frequency increases (Fig. 18 (b)), suggesting stronger attenuation and smoothing effects. Hence, bilinear upsampling typically biases the network towards generating smoother outputs, as prevalent in generative models [40]. Transposed convolutions, however, are not guaranteed to be low-passed as they are learnable. Due to the spectral bias of network layers, they may be low-passed during early training to still enable reconstruction, but the results could be unstable (green curves).
Convolutional layers. When the unlearnt upsampling operations are absent, ConvDecoder ($3\times 3$) still enables reconstruction while Deep Decoder ($1\times 1$) fails completely (Fig. 18(a)). A similar phenomenon is also reported in image denoising [5, 30]. This again be attributed to CNN’s inherent spectral bias and suggests that the size of the kernel also matters, further corroborated in Tab. 7.
Discussion. Results of this pilot experiment suggest that the spatial kernels with low-pass characteristics, either learnable or unlearnt, are crucial to the success of untrained network priors. In particular, bilinear upsampling with a fixed low-pass filter produces more stable and better results (blue curves).
D.2 Depth, Width and Skip Connections
Here, we demonstrate that insights gained about the unlearnt upsampling can aid in understanding the connection between architectural characteristics and the reconstruction task.
Experimental setup ii. For this large-scale validation, we experimented with an isotropic encoder-decoder architecture used in the original DIP, i.e., equal width and kernel size for all layers throughout the network. Design choices are detailed in Tab. 6. Experiments were performed on the publicly available 4×4$\times$4 × under-sampled multi-coil knee MRI from fastMRI database [24].
Table 6: Test bed for studying the architectural influences of an encoder-decoder untrained networks.
Archi. Type Depth (d) # of Skips (s) Width (w) Kernel Size (k) $\mathbf{A_{d}\mathunderscore_{s}\mathunderscore_{w}\mathunderscore_{k}}$2-L, 3-L, 4-L, 5-L, 8-L} {zero, half, full} {32, 64, 128, 256} $3\times 3,5\times 5$}
Table 7: Influences of typical architectural design choices in knee reconstruction. Deeper and/or Narrower architectures tend to perform better; skip connections influence the deep architectures more; larger kernels perform slightly better. $\mathbf{A_{8}\mathunderscore_{full}\mathunderscore_{32}\mathunderscore_{3}}$ performs the best (in lime); $\mathbf{A_{2}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ performs the worst (in red).
Width ($\downarrow$) Depth ($\uparrow$) Archi. PSNR SSIM Archi. PSNR SSIM Archi. PSNR SSIM Archi. PSNR SSIM $\mathbf{A_{2}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ 26.67 0.530 $\mathbf{A_{2}\mathunderscore_{full}\mathunderscore_{128}\mathunderscore_{3}}$ 27.12 0.543 $\mathbf{A_{2}\mathunderscore_{full}\mathunderscore_{64}\mathunderscore_{3}}$ 27.70 0.583 $\mathbf{A_{2}\mathunderscore_{full}\mathunderscore_{32}\mathunderscore_{3}}$ 28.47 0.641 $\mathbf{A_{3}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ 28.22 0.590 $\mathbf{A_{3}\mathunderscore_{full}\mathunderscore_{128}\mathunderscore_{3}}$ 28.59 0.605 $\mathbf{A_{3}\mathunderscore_{full}\mathunderscore_{64}\mathunderscore_{3}}$ 28.55 0.616 $\mathbf{A_{3}\mathunderscore_{full}\mathunderscore_{32}\mathunderscore_{3}}$ 29.25 0.660 $\mathbf{A_{4}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ 28.68 0.617 $\mathbf{A_{4}\mathunderscore_{full}\mathunderscore_{128}\mathunderscore_{3}}$ 28.95 0.622 $\mathbf{A_{4}\mathunderscore_{full}\mathunderscore_{64}\mathunderscore_{3}}$ 28.87 0.624 $\mathbf{A_{4}\mathunderscore_{full}\mathunderscore_{32}\mathunderscore_{3}}$ 29.70 0.671 $\mathbf{A_{5}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ 28.61 0.613 $\mathbf{A_{5}\mathunderscore_{full}\mathunderscore_{128}\mathunderscore_{3}}$ 28.87 0.615 $\mathbf{A_{5}\mathunderscore_{full}\mathunderscore_{64}\mathunderscore_{3}}$ 29.33 0.648 $\mathbf{A_{5}\mathunderscore_{full}\mathunderscore_{32}\mathunderscore_{3}}$ 29.81 0.680 $\mathbf{A_{8}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ 28.98 0.625 $\mathbf{A_{8}\mathunderscore_{full}\mathunderscore_{128}\mathunderscore_{3}}$ 29.33 0.637 $\mathbf{A_{8}\mathunderscore_{full}\mathunderscore_{64}\mathunderscore_{3}}$ 29.45 0.651 $\mathbf{A_{8}\mathunderscore_{full}\mathunderscore_{32}\mathunderscore_{3}}$ 30.04 0.695 Skip Connections (--) Kernel Size ($\uparrow$) $\mathbf{A_{2}\mathunderscore_{half}\mathunderscore_{256}\mathunderscore_{3}}$ 26.91 0.535 $\mathbf{A_{2}\mathunderscore_{zero}\mathunderscore_{256}\mathunderscore_{3}}$ 26.83 0.535 $\mathbf{A_{2}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ 26.67 0.530 $\mathbf{A_{2}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{5}}$ 26.98 0.550 $\mathbf{A_{4}\mathunderscore_{half}\mathunderscore_{256}\mathunderscore_{3}}$ 28.55 0.621 $\mathbf{A_{4}\mathunderscore_{zero}\mathunderscore_{256}\mathunderscore_{3}}$ 27.54 0.697 $\mathbf{A_{5}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ 28.61 0.613 $\mathbf{A_{5}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{5}}$ 28.82 0.624 $\mathbf{A_{8}\mathunderscore_{half}\mathunderscore_{256}\mathunderscore_{3}}$ 29.12 0.669 $\mathbf{A_{8}\mathunderscore_{zero}\mathunderscore_{256}\mathunderscore_{3}}$ 28.51 0.609 $\mathbf{A_{8}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ 28.98 0.625 $\mathbf{A_{8}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{5}}$ 29.12 0.634
Figure 19: Generalizability of different architectures on the masked regions.
What do deeper and narrower architectures produce? (Tab. 7). Theoretically, as the number of layers (depth) or channels (width) increases, the ability of the network to learn arbitrarily high frequencies (details, noise) is typically increased [38]. While this is true for width, we have found that the effect on depth turns out to be attenuated by unlearnt upsampling. As evidenced in Fig. 19, deeper architectures typically generate smoother images, exhibiting a stronger preference for low-frequency information, whereas shallower counterparts, even though they have fewer parameters, are more susceptible to noise and overfitting (red arrows). This is more evident when comparing the same architectures with just different upsamplers, where the architectures with bilinear upsampling (stronger attenuation) are less prone to overfitting than the ones using nearest neighbor (NN) upsampling (cyan vs. blue). Hence, it is not merely the number of parameters but the architectural characteristics promoting low frequencies that seem to be the primary reason for the high performance. Note that all these results are only achievable when unlearnt upsampling is involved (gray dashed curves).
Skip connections. Deep architectures with zero skip connection converge more slowly and may lead to over-smoothing as shown in Fig. 19 (red curves). Skip connections greatly alleviate this issue and introduce more details (cyan curves), which we speculate could be due to the "reduced effective up-sampling rate". Yet, excessive skip connections make a deep architecture behave similarly as a shallower one, generating more noise (Fig. 19 right). Overall, they exert a greater influence on deeper architectures ($\mathbf{A_{8}\mathunderscore_{zero}}<\mathbf{A_{8}\mathunderscore_{full}}<\mathbf{A_{8}\mathunderscore_{half}}$) compared to shallower ones ($\mathbf{A_{2}\mathunderscore_{full}}\approx\mathbf{A_{2}\mathunderscore_{zero}}$).
D.3 Validations on image denoising and inpainting.
We reaffirmed our observations above on image denoising and inpainting, as shown in Fig. 20, Fig. 22, Fig. 23 and Fig. 21.
Figure 20: Denoising experiments. (Left) In non-convolutional networks, removing the upsampling hampers the denoising capability, which cannot be compensated by merely adjusting the network to be more under-parameterized. Transposed convolutions result in a more rapid decline in performance than bilinear upsampling. (Right) Convolutional layers alone exhibit certain denoising effects but necessitate early stopping. The showcased image is from the classic dataset Set9 [8].
Figure 21: Deep architectures with zero skip connection converge more slowly, i.e., $\mathbf{A_{8}\mathunderscore_{zero}\mathunderscore_{256}}$
We argue that the understanding of the upsampling and its interactions with other architectural elements can help explain why deeper networks with fewer skip connections converge more slowly, generate smoother outputs and are less prone to overfitting. Concretely, the upsampling operation inserted in-between the decoder layer slows down the generation of high frequencies required for transforming the lower-resolution feature maps into the higher-resolution target image, primarily due to its role as a fixed low-pass filter. As the network depth increases, the degree of smoothness increases (Fig. 22). Skip connections notably accelerate the convergence (Fig. 21) and ameliorate the over-smoothing issue, likely due to the reduced "effective" upsampling rate. All these observations are consistent with our MRI experiments in Sec. 7.
Figure 22: Denoising experiments. Deeper architectures with few or no skip connections tend to generate smoother outputs compared to the shallower ones.
Figure 23: Inpainting experiments. Deeper architectures with few or no skip connections tend to generate smoother predictions for the masked regions than the shallower architectures. Skip connections make deep architectures perform similarly as the shallower ones.
E Comparisons with ZS-SSL-UNet
In the main paper, we have included the results of the ResNet version of ZS-SSL [48]. Here, we constructed a UNet variant of it, dubbed ZS-SSL-UNet. As shown in Tab. 8, the architecture type impacts not only DIP but also deep unrolling networks, and potentially a broader area, which worth future investigations.
Table 8: Quantitative evaluations. Runtime: mean $\pm$ std mins per slice.
[…]

### v3

@@
-Towards Architecture-Agnostic Untrained Network Priors for Image Reconstruction with Frequency Regularization
+Architecture-Agnostic Untrained Network Priors for Image Reconstruction with Frequency Regularization
 Abstract
-Untrained networks inspired by deep image prior have shown promising capabilities in recovering a high-quality image from noisy or partial measurements, without requiring training data.
-Their success has been widely attributed to the spectral bias acting as an implicit regularization induced by suitable network architectures.
-However, applications of such network-based priors often entail superfluous architectural decisions, overfitting risks, and slow optimization, all of which hinder their practicality.
-In this work, we propose efficient, architecture-agnostic methods for a more direct frequency control over the network priors: 1) constraining the bandwidth of the white-noise input, 2) controlling the bandwidth of the interpolation-based upsamplers, and 3) regularizing the Lipschitz constants of the layers.
-We show that even with just one extra line of code, the overfitting issues in underperforming architectures can be alleviated such that their performance gaps with the high-performing counterparts can be largely closed despite their distinct configurations, mitigating the need for architecture tuning.
-This then makes it possible to employ a more compact model to achieve similar or superior performance to larger models with greater efficiency.
-Our regularized network priors compare favorably with current supervised and self-supervised methods on MRI reconstruction and image inpainting tasks, serving as a stronger zero-shot baseline reconstructor.
-Our code will be made publicly available.
+Untrained networks inspired by deep image priors have shown promising capabilities in recovering high-quality images from noisy or partial measurements without requiring training sets.
+Their success is widely attributed to implicit regularization due to the spectral bias of suitable network architectures.
+However, the application of such network-based priors often entails superfluous architectural decisions, risks of overfitting, and lengthy optimization processes, all of which hinder their practicality.
+To address these challenges, we propose efficient architecture-agnostic techniques to directly modulate the spectral bias of network priors: 1) bandwidth-constrained input, 2) bandwidth-controllable upsamplers, and 3) Lipschitz-regularized convolutional layers.
+We show that, with just a few lines of code, we can reduce overfitting in underperforming architectures and close performance gaps with high-performing counterparts, minimizing the need for extensive architecture tuning.
+This makes it possible to employ a more compact model to achieve performance similar or superior to larger models while reducing runtime.
+Demonstrated on inpainting-like MRI reconstruction task, our results signify for the first time that architectural biases, overfitting, and runtime issues of untrained network priors can be simultaneously addressed without architectural modifications.
+Our code is publicly available  [^1] https://github.com/YilinLiu97/Untrained-Recon.git.
 1 Introduction
 Magnetic resonance imaging (MRI) is a mainstream imaging tool for medical diagnosis.
-Reconstructing MR images from raw measurements refers to the transformation from Fourier spectrum of the object in $k$-space to image space.
+Reconstructing MR images from raw measurements involves data transformation from a Fourier spectrum in $k$-space to image space [11, 28].
 Since acquiring full $k$-space measurements is time-consuming, under-sampled $k$-space data are often collected to reduce scan times.
-Accelerated MRI is thus known as an ill-posed inverse problem that conventionally requires handcrafted priors [31, 27] to mitigate the resulting aliasing artifacts in the output images.
-While supervised learning methods based on convolutional neural networks (CNNs) have demonstrated better reconstruction quality with fewer measurements, their training relies on paired under-sampled and fully-sampled measurements, which are expensive to acquire and could raise issues on robustness and generalization when the acquisition protocol or anatomy changes [23, 24].
-Figure 1: Example results from underperforming architectures with 4×4$\times$4 × under-sampling.
-Turning the left to the right simply by low-pass filtering the white-noise input via a Gaussian blur kernel, which can be implemented as few as one or two lines of code.
-SSIM (normal-$\uparrow$) values are reported.
-Instead of requiring large-scale datasets for capturing prior statistics, the untrained network inspired by deep image prior (DIP) [45] requires only the corrupted or partial measurements and regularizes the reconstruction solely through the inductive biases of its architecture, making it a representative zero-shot self-supervised reconstruction method for a wide range of natural and medical applications [50, 28, 37].
-Concretely, DIP parameterizes the unknown desired image via a neural network and optimizes the network parameters such that the output image transformed by the degradation model matches the acquired measurements.
-Such parameterization is observed to exhibit high resistance to noise and corruption, acting as a form of implicit regularization.
-Studies have attributed this property to CNN’s inherent spectral bias – the tendency to fit the low-frequency signals before the high-frequency signals (e.g., noise) [41, 5], where the choice of network architecture is often shown to be critically relevant [5, 30, 1].
-Despite its great promise, favorable results obtained from the untrained network prior are contingent upon two critical factors: an optimal architecture for a specific task and an early-stopping strategy to prevent overfitting to the noisy, partial measurements.
-Furthermore, while being agnostic to domain shifts by optimizing on a per-image basis [48], such reconstruction process is inherently slow, hindering its practicality.
-While these issues are often intertwined — inappropriate and over-parameterized architectures usually exacerbate overfitting and increase runtime, most prior works tackle these challenges separately.
+This makes accelerated MRI reconstruction an ill-posed inverse problem that conventionally requires handcrafted priors [30, 25] to mitigate the resulting aliasing artifacts.
+While supervised learning methods based on convolutional neural networks (CNNs) have enhanced reconstruction quality with fewer measurements, their training relies on paired under-sampled and fully-sampled measurements, which are expensive to acquire and can affect robustness and generalization across different acquisition protocols or anatomical variations [21, 22].
+Figure 1: Improving underperforming architectures (SSIM ($\uparrow$)).
+Turning the left to the right simply by replacing the white-noise input with selected Fourier features or low-pass filtering the noise input, which can be implemented with a few lines of code.
+Instead of requiring large-scale datasets for capturing prior statistics, untrained networks [12] inspired by deep image prior (DIP) [43] use only the corrupted or partial measurements and a task-specific forward operator.
+Reconstruction is regularized solely by the inductive biases of the network architectures, enabling zero-shot self-supervised reconstruction for various imaging inverse problems [48, 26, 36].
+Concretely, a CNN, which parameterizes the unknown desired image, is optimized such that the output image, transformed by the forward operator, matches the acquired measurements.
+Such parameterization exhibits surprisingly high resistance to noise and corruption, which acts as a form of implicit regularization.
+Recent studies have attributed this property to CNN’s inherent spectral bias—the tendency to fit the low-frequency signals before the high-frequency signals (e.g., noise) [39, 4].
+The choice of network architecture is shown to be critically relevant to such bias [4, 29, 1].
+Despite the great promise, obtaining favorable results with untrained network priors is contingent upon two critical factors: an optimal architecture specific to a task and an early-stopping strategy to prevent overfitting to noisy or partial measurements.
+Furthermore, optimizing on a per-image basis makes the reconstruction process domain-agnostic but inherently slow [46].
+While these issues are intertwined, with overfitting and runtime issues exacerbated by inappropriate and over-parameterized architectures, most existing efforts tackle these challenges separately.
 For architectural design, existing methods rely on either handcrafting or utilizing neural architecture search techniques [12, 6, 1, 5, 24, 14].
-However, the lack of a consensus on the desired architectural prior often leads to a laborious search.
-Another line of work is dedicated to preventing overfitting through early-stopping [48, 46], subspace optimization [2] or pretraining-then-fine-tuning [3, 35], which also helps shorten the runtime.
-Similar in spirit to traditional ways of combating overfitting, these methods aim to use fewer trainable parameters with little performance degradation (and may involve costly pretraining) or a held-out subset of measurements for self-validation.
-In this work, we explore the possibility of modulating the frequency bias and hence the regularization effects of network priors in an architecture-agnostic manner, aiming to enhance a given architecture irrespective of its specific configurations.
-This is conceivable, in light of the recent body of theoretical and empirical evidence indicating that there are only a few key componenets within the architecture that are driving forces behind the spectral bias in DIP [30, 5, 41, 15].
-Inspired by these findings, we develop simple yet effective methods from a frequency perspective to efficiently regularize the network priors, alleviating overfitting by curbing the over-fast convergence of high frequencies, all with minimal architectural modifications and computational costs.
-Specifically, we propose to (i) constrain the effective bandwidth of the white-noise input via simple spatial blurring or selected Fourier features, (ii) adjust the bandwidths of the interpolation-based upsamplers to control their attenuation (smoothing) extents, and (iii) regularize the Lipschitz constants of the layers to enforce function smoothness.
-We empirically find that our regularized network priors not only exhibit less vulnerability to overfitting but also tend to achieve better extrapolation and thus higher peak performance in inpainting-like tasks.
-Demonstrated in the context of MRI reconstruction, our methods significantly improve models across various architectural configurations, largely reducing the need for extensive architectural tuning (Fig.
+However, the lack of consensus on architectural priors often leads to laborious search.
+Another line of work is dedicated to preventing overfitting through oracle early-stopping criterions [46, 44], subspace optimization [2] or pretraining then fine-tuning [3, 34].
+These methods mostly use fewer trainable parameters or hold out a subset of measurements for self-validation, in the spirit of traditional strategies, and often trade-off accuracy or involve costly pre-training.
+In this work, we explore the possibility of modulating the frequency bias and hence the regularization effects of network priors in an architecture-agnostic manner, aiming to enhance the performance of a given architecture irrespective of its configuration specifics.
+This is conceivable in light of the recent body of theoretical and empirical evidence indicating that there are only a few key components (e.g., unlearnt upsampling) within the architecture that are the driving forces behind the spectral bias in DIP [29, 4, 39, 13].
+Motivated by these findings, we develop efficient methods from a frequency perspective to effectively regularize the network priors, alleviating overfitting by curbing the overly rapid convergence of undesired high-frequency components, all with minimal architectural modifications and computational costs.
+Specifically, we propose to (1) constrain the effective bandwidth of the input via blurring or using Fourier features with a narrower frequency range, (2) adjust the bandwidths of the interpolation-based upsamplers with controllable attenuation (smoothing) extents, and (3) regularize the Lipschitz constants of the convolutional layers to enforce function smoothness.
+We found empirically that by mitigating convergence to high-frequency components, our regularized network priors not only exhibit less vulnerability to overfitting but also tend to achieve better extrapolation capabilities in inpainting tasks.
+In the context of MRI reconstruction, which is essentially an inpainting task occurring in $k$-space, our methods significantly improve models across various architectural configurations without necessitating extensive architectural tuning (Fig.
 1).
-The efficacy of our methods is also showcased in image inpainting and denoising tasks, where the control over frequency is adjusted accordingly.
-By minimizing architectural influences, our approach additionally offers a unique advantage in efficiency: a smaller, previously underperforming network, can now achieve on-par performance or even surpass a larger, heavily parameterized high-performing network.
-Our contributions are three-fold:
-  • We propose efficient methods that directly modulate the frequency bias in the untrained network priors in an architecture-agnostic manner, effectively addressing challenges related to architecture, overfitting, and runtime in a unified approach.
-  • The enhanced untrained networks match the performance of leading self-supervised methods and surpass the supervised method especially on out-of-domain data, with up to 86×86$\times$86 × reduction in runtime ($\sim 5$ mins/slice) for MRI reconstruction, serving as a stronger zero-shot reconstruction baseline than the original DIP scheme.
-  • Our empirical findings and results on medical and natural image reconstruction tasks shed light on the spectral behaviors of network-based priors—CNNs operating in a single-instance generative setting.
+Their efficacy is also showcased in denoising and inpainting for natural images.
+By minimizing architectural influences, our approach additionally offers a unique advantage in efficiency: a smaller, previously underperforming network, can now achieve performance on-par with or even surpasses a larger, heavily parameterized high-performing network.
+Our contribution is three-fold:
+  • We propose efficient techniques to directly modulate the frequency bias in untrained network priors, addressing architectural design, overfitting, and runtime challenges in a unified, architecture-agnostic manner.
+  • The enhanced untrained networks match leading self-supervised methods with up to 90$\times$ faster runtime (1 hr/slice to $\sim$5  min/slice) for MRI reconstruction and surpass supervised methods on out-of-domain data.
+  • Our findings on medical and natural image reconstruction reveal the spectral behaviors of CNNs in a single-instance generative setting.
 2 Related Work
-Function Smoothness and Spectral Bias.
+Spectral Bias and function smoothness.
 Function smoothness, also referred to as function frequency, quantifies how much the output of a function varies with changes in its input [9].
 Spectral bias [37, 45] is an implicit bias that favors learning functions changing at a slow rate (low-frequency), e.g., functions with a small Lipschitz constant.
-In visual domains, this is evident in the network’s output lacking subtle details.
-Many regularization techniques shown to aid generalization encourage smoothness implicitly, such as early stopping, $\ell_{2}$ regularizer [39].
-Smoothness has been widely used as a model complexity measure in place of model size to account for the well-known "double-descent" phenomenon associated with generalization [34].
+In visual domains, this is evident from the lack of subtle details in network outputs.
+Many regularization techniques to aid generalization, such as early stopping and $\ell_{2}$ regularization [38, 33], implicitly encourage smoothness.
 To explicitly promote smoothness, it is natural to penalize the norm of the input-output Jacobian [35, 15].
-However, due to the high dimensionality of the output such as in accelerated MRI, computation of the Jacobian matrix during training is often intractable.
+However, computation of the Jacobian matrix for the high-dimensional MRI data during training is very expensive.
 Another efficient and prevalent solution is to constrain the network to be c-Lipschitz with a pre-defined Lipschitz constant c [32, 10].
-We develop a suitable form of Lipschitz regularization for untrained networks by instead penalizing learned Lipschitz constants, with a novel aim of achieving architecture-agnostic untrained image reconstruction.
-Input Frequency and Generalization.
-Input has played an important role in helping neural networks represent signals of various frequencies.
-As in implicit neural representations (INRs) [42] and neural radiance fields (NeRFs) [32] where coordinates are mapped to RGB values, naively training with raw coordinates as inputs results in over-smoothing; encoding the input coordinates with sinusoidal functions of higher frequencies enables the network to represent higher frequencies [32, 43].
-Rahaman et al.
-[38] also show theoretically and empirically that fitting becomes easier for the network when the input itself contains high-frequency components.
-However, it has recently been reported that the high-frequency positional input encodings lead to the failure of NeRFs in few-shot settings due to over-fitting [49].
-Here, we show that similar issue also applies to untrained network priors and can be efficiently addressed by our methods.
-Avoid Overfitting in Untrained Networks.
-A line of effort has been exclusively devoted to preventing overfitting to noisy images or measurements.
+We followed this line with a novel aim of achieving architecture-agnostic untrained image reconstruction.
+Input frequency and overfitting.
+The network input plays an important role in helping the neural network represent signals of various frequencies.
+Pioneering work on spectral bias [37] showed theoretically and empirically that fitting high frequencies becomes easier, provided that the data manifold itself contains high-frequency components (Sec.
+4 in [37]).
+This has directly motivated implicit neural representations (INRs) [40] and neural radiance fields (NeRFs) [31] where coordinates are mapped to RGB values: naively training with raw coordinates as inputs results in over-smoothing; encoding the input coordinates with sinusoidal functions of higher frequencies enables the network to represent higher frequencies [31, 41].
+However, it has recently been reported that the high-frequency bands of NeRF’s input encodings incur overfitting and lead to failure in few-shot settings [47].
+We found similar issues in untrained networks and propose to constrain the input’s frequency range to counteract the convergence of high frequencies.
+Architecture-induced spectral bias.
+Recent studies on the working mechanisms of DIP reveal that unlearnt upsampling, with the low-pass characteristics of its interpolation filter, is responsible for the regularizing effects of DIP [29, 13, 4].
+Liu et al.
+[29] showed that the fixed upsampling operations readily bias the architecture towards producing low-frequency outputs, critically influencing both the peak PSNR and the starting point at which performance decays.
+The convolutional layer is another core element that exhibits stronger frequency selectivity compared to fully-connected layers and 1D layers [4, 39].
+These findings motivated us to operate directly on these elements to achieve architecture-agnostic control.
+Avoiding overfitting is a primary goal of unsupervised reconstruction where only noisy or partial measurements are available.
 Wang et al.
-[46] propose to track the running variance of the output for an early-stopping criterion, but it is found to be unstable in medical image reconstruction [2].
+[44] proposed an early-stopping (ES) criterion by tracking the running variance of the output, but it is found to be unstable in medical settings [2].
 Yaman et al.
-[48] propose to split the available measurements into a training and a validation subset and use the latter for self-validation for automated early stopping.
-Transfer-learning-based untrained networks perform pre-training on synthetic data followed by fine-tuning [3, 35] or subspace optimization [2].
-These methods aim to use fewer trainable parameters to avoid overfitting with little performance degradation.
-In contrast, our methods alleviate overfitting from a frequency perspective and enable significantly better performance while maintaining the same or less model-wise and computation-wise complexity.
+[46] proposed to split the available measurements into training and validation subsets and use the latter for self-validation for automated early stopping.
+While ES prevents overfitting, it cannot enhance the capability of a given architecture like our approach.
+Transfer-learning based methods aim to use fewer non-trainable parameters by performing pre-training followed by fine-tuning [3, 34] or subspace optimization [2].
+In contrast, our method directly modulates the spectral bias to mitigate the convergence of undesired high frequencies, which is found to also improve the networks’ extrapolation capabilities, all while maintaining the model complexity.
 Figure 2: Overview of the proposed regularized network priors for MRI reconstruction.
+By adjusting the bandwidth of the input, the interpolation-based upsampling, and regularizing the convolutional layers, our approach enables more direct control over the spectral bias of architectures with various depths and widths.
 3 Method
 3.1 Preliminaries
-Accelerated MRI The goal of accelerated MRI reconstruction is to recover a desired image $\mathbf{x\in\mathbb{C}}^{n}$ ($n=n_{h}\times n_{w}$) from a set of under-sampled k-space measurements.
-We focus on a multi-coil scheme in which the forward model is defined as:
-$\mathbf{y}_{i}=\mathbf{A}_{i}\mathbf{x}+\epsilon\text{ with }\mathbf{A}_{i}=\mathbf{MFS}_{i},\quad i=1,\dots,c,$ (1)
+Accelerated MRI The goal of accelerated MRI reconstruction is to recover a desired image $\mathbf{x\in\mathbb{C}}^{n}$ ($n=n_{\text{h}}\times n_{\text{w}}$) from a set of under-sampled k-space measurements.
+We focus on a multi-coil scheme in which the forward model is defined as
+$\mathbf{y}_{i}=\mathbf{A}_{i}\mathbf{x}+\epsilon,\quad\mathbf{A}_{i}=\mathbf{MFS}_{i},\quad i=1,\dots,c,$ (1)
 where $\mathbf{y}_{i}\in\mathbb{C}^{m}$ denotes the $k$-space measurements from coil $i$, $c$ denotes the number of coils, $\mathbf{S}_{i}\in\mathbb{C}^{n}$ denotes the coil sensitivity map (CSM) that is applied to the image $\mathbf{x}$ through element-wise multiplications, $\mathbf{F}\in\mathbb{C}^{n\times n}$ denotes the 2D discrete Fourier transform, $\mathbf{M}\in\mathbb{C}^{m\times n}$ denotes the under-sampling mask, and $\epsilon\in\mathbb{C}^{m}$ denotes the measurement noise.
@@
 The image $\mathbf{x}$ is parameterized via a neural network $\mathbf{G_{\theta}(z)}$ with a fixed noise input vector $\mathbf{z}$ drawn from a uniform distribution $z\sim\mathcal{U}(0,1)$.
-With the aforementioned MRI forward model, the untrained network solves the following optimization problem:
+With the MRI forward model in Eq.
+1, the untrained network solves the following optimization problem:
 $\theta^{\ast}=\mathop{\arg\min}\limits_{\theta}\mathcal{L}(\mathbf{y};\mathbf{AG_{\theta}(z)}),\quad\mathbf{x}^{\ast}=\mathbf{G_{\theta^{\ast}}(z)}.$ (2)
-Such parameterization allows novel image priors to be designed dependent on the network architecture and the associated parameters.
-Nevertheless, many studies augment the untrained networks with traditional image regularizers [29], e.g., Total variation (TV), though it can only partially alleviate over-fitting [35, 2].
-In our experimental sections, we show that TV is not as effective as ours in improving under-performing architectures.
+This parameterization enables the design of novel image priors based on the network architecture and its parameters, rather than on handcrafted image space priors.
+Nevertheless, many studies augment the untrained networks with traditional image regularizers [27], e.g., total variation (TV), which promotes piecewise constant images and can only partially alleviate over-fitting [34, 2].
 3.2 Architecture-Agnostic Frequency Regularization
-To modulate the regularization effects of the network prior, we identify three sources of spectral bias within the network and introduce corresponding regulation methods, as depicted in Fig.
+To modulate the regularization effects of the network prior, we identify three core elements within the framework of DIP that lead to spectral bias and introduce corresponding regulation methods, as depicted in Fig.
 2.
-Each method requires only minimal hyperparameters to achieve architecture-agnostic control over the network priors.
 Figure 3: (a) Visualization of the 1D white noise, the low-passed noise via Gaussian blur, and Fourier features of various frequencies in the frequency domain.
-(b) Gaussian blurred noise input and Fourier features with lower $f_{c}$ ($L=4$ or $8$) are both effective in alleviating overfitting and enhancing the peak performance.
+(b) Limiting the input’s bandwidth via either Gaussian blur or Fourier features with lower $f_{c}$ ($L=4$ or $8$) are both effective in alleviating overfitting and enhancing the peak performance.
 Bandwidth-Constrained Input.
-An important aspect of untrained networks that can be easily overlooked is their inputs.
-Conventionally, the inputs are randomly sampled from either a uniform or Gaussian distribution and are then mapped to an image.
-From a frequency perspective, such white-noise input comprises all frequencies with uniform intensities [11].
+Inspired by implicit neural representations (INRs), we rethink the role of inputs in untrained networks for learning different frequencies.
+Conventionally, the inputs are randomly sampled from either a uniform or Gaussian distribution and are then mapped to image intensities.
+From a frequency perspective, such input comprises all frequencies with uniform intensities [8], as white noise with variance $\sigma^{2}$ exhibits an autocorrelation that is a scaled Dirac $\delta$-function $\sigma^{2}\delta(t)$, whose Fourier transform $\mathcal{F}$ has a constant magnitude $\sigma^{2}$ spanning all frequencies $\mu$, i.e., $\mathcal{F}\{\sigma^{2}\delta(t)\}(\mu)=\sigma^{2}$, $\mu\in\mathbb{R}$.
 With this view in mind, we draw an analogy between untrained networks and INRs that map Fourier features to RGB values.
-Fourier features are sinusoid functions of the input coordinates $\mathrm{\mathbf{p}}$, i.e., $[\sin(\mathrm{2^{0}\pi\mathbf{p}}),\cos(\mathrm{2^{0}\pi\mathbf{p}}),...,\sin(2^{L-1}\pi\mathbf{\mathrm{\mathbf{p}}}),\cos(2^{L-1}\pi\mathbf{\mathrm{\mathbf{p}}})]$, where a larger $L$ assists the network in representing higher-frequency functions [43].
-Figure 4: Narrowing input’s bandwidth enhances performance as a "free lunch". The output becomes smoother as $\sigma$ increases, up to a certain point.
-White noise with variance $\sigma^{2}$ exhibits an autocorrelation that is a scaled Dirac $\delta$-function $\sigma^{2}\delta(t)$, whose Fourier transform $\mathcal{F}$ has a constant magnitude $\sigma^{2}$ spanning all frequencies $\mu$, i.e., $\mathcal{F}\{\sigma^{2}\delta(t)\}(\mu)=\sigma^{2}$, $\mu\in\mathbb{R}$.
+Fourier features are sinusoid functions of the input coordinates $\mathrm{\mathbf{p}}$, i.e., $[\sin(\mathrm{2^{0}\pi\mathbf{p}}),\cos(\mathrm{2^{0}\pi\mathbf{p}}),...,\sin(2^{L-1}\pi\mathbf{\mathrm{\mathbf{p}}}),\cos(2^{L-1}\pi\mathbf{\mathrm{\mathbf{p}}})]$, where a larger $L$ assists the network in representing higher frequencies [41].
+This is consistent with [37] which indicates that the frequency magnitudes that can be expressed by the network increase with increasing frequency in the data manifold.
 In this sense, an untrained network can be viewed as mapping a broad spectrum of Fourier features to a target image (Fig.
 3 (a)).
-We hypothesize that this enhances the network’s representation ability but likely incurs over-fitting due to the faster convergence of high-frequency components.
-To validate this, we applied a Gaussian blur filter $\mathcal{G}_{s,\sigma}$ on the white-noise input $z$ to remove a certain amount of high frequencies before passing it to the network, defined as $z*\mathcal{G}_{s,\sigma}$, where *** denotes convolution.
+We hypothesize that this promotes rapid convergence of high frequencies, which is likely to bias the network towards high-frequency artifacts and hinder its ability to exploit
+Figure 4: Narrowing input’s bandwidth promotes low frequencies and enhances extrapolation capabilities as a "free lunch". The output becomes smoother as $\sigma$ increases, up to a certain point.
+spatial information effectively for extrapolation.
+To validate this, we applied a Gaussian blur filter $\mathcal{G}_{s,\sigma}$ on the noise input $z$ to remove a certain amount of high frequencies before passing it to the network: $z*\mathcal{G}_{s,\sigma}$, where * denotes convolution.
 The filter size $s$ and sigma value $\sigma$ that controls the filter’s bandwidth are hyperparameters.
@@
 4, simply adjusting $\sigma$ already brings significant gains without architectural changes.
-Similarly, substituting the noise input with Fourier features, with a carefully selected maximum frequency $f_{c}\propto L$ to limit the input’s effective bandwidth has also proven effective in Fig.
-3 (b) and our ablation study (Tab.
-23), further validating our hypothesis.
-Fourier features introduce frequency-diverse input akin to white noise but in a controlled manner, enabling frequency regularization.
-Drawing from these empirical findings, we propose to regularize the input’s frequency range, achievable via simple spatial blurring or selected Fourier features, to efficiently mitigate overfitting and enhance network performance.
-While Gaussian blur and Fourier Features both offer bandwidth control, they yield different frequency representations (e.g., Fourier features are more discrete than Gaussian blurred noise), which may affect performance and susceptibility to parameter tuning differently.
-We present and compare both results in the experiments.
-We note that inputs have been deemed crucial in guiding the generation of the final image not only in INRs but also in generative networks, e.g., transformed Fourier features have been used as an alternative to random input in StyleGAN3 [21], for a different goal of improving the architecture’s equivariance.
-In contrast, we use a low-passed version of Fourier features or noise input to prevent over-fast convergence on high frequencies and enable better extrapolation.
-Similar in spirit, recent work by [49] also shows that masking the high-frequency Fourier features helps NeRFs generalize in few-shot settings.
+Alternatively, as shown in Fig.
+3 (b), substituting the noise input with Fourier features, with a carefully selected maximum frequency $f_{c}\propto L$ (e.g., $L=4$ or 8) to narrow the input’s effective bandwidth, is also effective.
+As $L$ increases, e.g., $L=16$, the frequency range of Fourier features input approximates that of the original noise input, and the performance deteriorates similarly, further supporting our hypothesis.
+Fourier features introduce frequency-diverse input akin to white noise but in a controlled manner, enabling regularization over the frequency selectivity of the network.
+In light of this empirical evidence, we propose to limit the input’s bandwidth to mitigate the fitting of high-frequency components in untrained networks, achievable efficiently via blurring or Fourier features with a narrowed frequency range.
+Gaussian blur and Fourier Features offer flexible bandwidth control over the input through hyperparameters $\{s,\sigma\}$ and $L$, respectively, which allows for scaling under higher noise level/undersampling rates (examples of 8×8$\times$8 × undersampling in Fig.
+8).
+We experimented with more sophisticated schedulers for the hyperparameters, but found that simply fixing them throughout the training yields superior performance.
 Table 1: Influences of upsamplers on reconstruction.
@@
 (Millions) ConvDecoder [6] 28.69 $\pm$ 1.6 31.78 $\pm$ 1.2 32.31 $\pm$ 1.3 32.48 $\pm$ 1.2 4.1 M Deep Decoder [12] 24.55 $\pm$ 1.1 27.10 $\pm$ 0.9 31.36 $\pm$ 1.4 32.68 $\pm$ 1.1 0.47 M
-Smoothness-controlled Upsampling.
-We found that constraining the input’s bandwidth significantly enhances shallower architectures, yet the improvement diminishes as the network depth increases.
-This could partly be attributable to the fact that subsequent layers along with nonlinearity can always generate new arbitrarily high frequencies, as theoretically and empirically proved in [38, 21].
-We notice that the interpolation-based upsampling methods within the network, such as nearest neighbor and bilinear, are essentially implicit low-pass filters smoothing out the alias frequencies caused by the increased sampling rate.
-Prior works [30, 15] on image denoising have shown that these non-trainable upsampling methods are driving forces behind the spectral bias of the network priors, delaying the convergence of higher frequencies.
-Different upsamplers bias the network towards different spectral properties [30].
-We confirm that upsampling also substantially influences image reconstruction and further reveal that an appropriate upsampling method can bridge the performance gap between two distinct architectures (Tab.
-1).
-Motivated by these results, we introduce a controllable upsampler to modulate the network’s spectral bias, especially the deeper architectures.
-We achieve this by first interleaving the input feature maps with zeros, and then convolving them with a customized low-pass filter with adjustable bandwidth (Fig.
+Bandwidth-Controllable Upsampling.
+We observed that only constraining the input’s bandwidth significantly enhances shallower architectures, yet the improvement diminishes as the network depth increases (Tab.
+23).
+This could be partly attributed to the increased network layers that can always generate new arbitrarily high frequencies [37, 19].
+We note that the interpolation-based upsampling methods within the network, such as nearest neighbor and bilinear, essentially act as implicit low-pass filters, attenuating the alias frequencies caused by the increased sampling rate of the input feature maps.
+Prior works [29, 13] in denoising have shown that these non-trainable upsampling methods are driving forces behind the spectral bias of DIP, delaying the convergence of higher frequencies.
+Different upsamplers bias the network towards different spectral properties, depending on the bandwidth of the interpolation filter [29].
+Here, we show in Tab.
+1 that upsampling also substantially influences the performance image reconstruction.
+Motivated by these results, we introduce an upsampler with controllable bandwidth so as to modulate the network’s spectral bias, especially for deeper architectures.
+We construct it by 1) first interleaving the input feature maps with zeros, and then 2) convolving them with a customized low-pass filter with adjustable bandwidth (Fig.
 2).
-For filter design, we adopt the Kaiser-Bessel window as it offers explicit control over the tradeoffs between passband ripple and stopband attenuation.
+For filter design, we adopt the Kaiser-Bessel window [18] as it offers explicit control over the tradeoffs between passband ripple and stopband attenuation.
 The Kaiser window is defined as
 $w(n)=I_{0}(\beta\sqrt{1-(2n/M)^{2}})/I_{0}(\beta),-M/2\leq n\leq M/2,$ (3)
-where $M$ is the desired spatial extent of the window, $\beta\geq 0$ is the shape parameter—the higher it is, the greater the stopband attenuation is (and generally the smoother the image is), and $I_{0}$ is the zeroth-order modified Bessel function of the first kind [20].
+where $M$ is the desired spatial extent of the window, $\beta\geq 0$ is the shape parameter—the higher it is, the greater the stopband attenuation is (and generally the smoother the image is), and $I_{0}$ is the zeroth-order modified Bessel function of the first kind.
 This plug-and-play upsampler can be inserted in different layers with different $M$ and $\beta$ hyperparameters, offering flexible and precise control.
 Lipschitz-Regularized Layers.
-Spectral bias towards low frequencies favors functions that do not change at a high rate, i.e., functions with small Lipschitz constants.
-A function $f:\mathcal{X}\rightarrow\mathcal{Y}$ is said to be Lipschitz continuous if there is a constant $k>0$ such that
-$\|f(\mathrm{x_{1}})-f(\mathrm{x_{2}})\|_{\text{p}}\leq k\|\mathrm{x_{1}}-\mathrm{x_{2}}\|_{\text{p}}\quad\forall\mathrm{x_{1}},\mathrm{x_{2}}\in\mathcal{X},$ (4)
-where $k$ is the Lipschitz constant that bounds how fast $f$ can change globally w.r.t. input perturbations.
+Compared to the non-trainable upsampling that only attenuates signals, the network layer with nonlinearies is the only operation capable of generating new frequencies [19].
+We regularize their Lipschitz constants to control their sensitivity to input variations, which in turn can affect the spectral bias.
+Formally, a function $f:\mathcal{X}\rightarrow\mathcal{Y}$ is said to be Lipschitz continuous if there is a constant $k>0$ such that $\|f(\mathrm{x_{1}})-f(\mathrm{x_{2}})\|_{\text{p}}\leq k\|\mathrm{x_{1}}-\mathrm{x_{2}}\|_{\text{p}}\quad\forall\mathrm{x_{1}},\mathrm{x_{2}}\in\mathcal{X},$ where $k$ is the Lipschitz constant that bounds how fast $f$ can change globally w.r.t.
+input perturbations.
+Spectral bias towards low frequencies favors functions with small Lipchitz constants.
 Instead of upper bounding the Lipschitz constants of the network layers to pre-defined and manually chosen values as in [39], we make the per-layer Lipschitz bounds learnable and regularize their magnitudes during optimization.
 The Lipschitz constant of a convolutional layer is bounded by the operator norm of its weight matrix [10].
-To bound a convolutional layer to a specific Lipschitz constant $k$, the layer with $m$ input channels, $c$ output channels and kernels of size $w\times h$ is first reshaped to a 2-D matrix $W^{\ell}\in\mathbb{R}^{n\times cwh}$, and then normalized as:
+To bound a convolutional layer to a specific Lipschitz constant $k$, the layer with $m$ input channels, $c$ output channels and kernels of size $w\times h$ is first reshaped to a 2-D matrix $W^{\ell}\in\mathbb{R}^{n\times cwh}$, and normalized as
 $\tilde{W_{\ell}}=\frac{W_{\ell}}{\text{max}(1,\frac{\|W_{\ell}\|_{\text{p}}}{\mathtt{SoftPlus}(k_{\ell})})},$ (4)
@@
 Such formulation only normalizes $W_{\ell}$ if its matrix norm is larger than the learned Lipschitz constraint during training.
-Integrating the ultimate Lipschitz regularization with learned Lipschitz constants into Eq.
-2, our regularized training objective is:
+Integrating the ultimate Lipschitz regularization into Eq.
+2, our regularized training objective is
 $\min_{\Theta,K}\mathcal{L}(\mathbf{y};\mathbf{AG_{\Theta}(z)})+\lambda\sum_{l=1}^{L}\mathtt{SoftPlus}(\mathbf{k}_{\ell})^{2}$ (5)
@@
 4.1 Setup and Datasets
-We first (1) validate the proposed methods’ effectiveness in enhancing the performance of untrained networks across various architectural configurations, especially those originally underperforming.
+We first (1) validate the effectiveness of the proposed methods in enhancing the performance of untrained networks across various architectural configurations, especially those originally underperforming.
 We then (2) benchmark the enhanced versions of those compact architectures against established supervised and self-supervised methods on both in-domain and out-of-domain datasets in terms of accuracy and efficiency.
 We also (3) compare our methods with self-validation-based early stopping [46] on overcoming overfitting, and show that it is complementary to our approach by further shortening the reconstruction time.
-Finally, we demonstrate the utility of our methods in (4) general image inpainting and denoising tasks and perform spectral bias analysis on all evaluated tasks.
-The MRI experiments were performed on two publicly available datasets including the multi-coil knee and brain MRI images from fastMRI database [25], and multi-coil knee MRI images from Stanford 3D FSE knee dataset [19].
+Finally, we demonstrate the utility of our methods in (4) general image inpainting and denoising tasks and (5) perform spectral bias analysis on all evaluated tasks.
+The MRI experiments were performed on two publicly available datasets: the multi-coil knee and brain MRI images from fastMRI database [23], and multi-coil knee MRI images from Stanford 3D FSE knee dataset [17].
 The fully-sampled $k$-space data was retrospectively masked by selecting 25 central $k$-space lines along with a uniform undersampling at outer $k$-space, achieving the standard 4×4$\times$4 × acceleration.
-For training the supervised baseline, the knee training set consists of 367 PD and PDFS slices and the brain training set consists of 651 slices with a mixture of T1, T2 weighted images.
+For training the supervised baseline, the knee training set consists of 367 PD and PDFS slices and the brain training set consists of 651 slices with a mixture of T1 and T2 weighted images.
 50 knee slices and 50 brain slices were sampled from the respective multi-coil validation datasets for evaluation.
@@
 $M$ and $\beta$ for the Kaiser-based upsamplers are chosen to be $\{15\times N-1,5\}$ and $\{5\times N\}$ for knee data ($N$ denotes the nth-level), and $\{5\times N\}$ and $\{5\times N\}$ for brain data, respectively.
-$\lambda$ is set to $1$ for Lipschitz regularizer.
+$\lambda$ is set to $1$ for the Lipschitz regularizer.
 Table 2: Effectiveness of the methods in bridging performance gaps among different architectures, evaluated on fastMRI brain datasets.
@@
 Fig.
-5 gives a quantitative overview of the substantial improvement using our approach in architectures with diverse configurations on both knee and brain datasets.
+5 gives a quantitative overview of the substantial improvement give by our approach in architectures with diverse configurations on both knee and brain datasets.
 The different results of the original architectures also confirm the influences of architectural choices on performance.
@@
 4.5 and appendix, these underperforming architectures learn high frequencies more quickly (though this may be desired for other tasks [29]) and are more susceptible to overfitting, incurring severe artifacts in the output (Figs.
-6 and 9).
+6 and 10).
 When applied with our methods, as detailed in Tab.
 2 and Tab.
-3, a large boost in performance is observed in all architectures, especially in $\mathbf{A_{2}\mathunderscore_{256}}$.
+3, a large boost in performance is observed in all architectures, especially $\mathbf{A_{2}\mathunderscore_{256}}$.
 We observe that using low-passed inputs via either selected Fourier features or blurring brings the most benefits to the shallower architectures.
@@
 4.2), with the knee data requiring greater attenuation.
-This is also consistent with previous findings that they require different numbers of channels for the architecture [9].
+This is also consistent with previous findings that they require different numbers of channels for the architectural choices [6].
 Our methods greatly alleviate the need for such architectural tuning by instead allowing for the adjustment of a few key hyperparameters.
@@
 Figure 7: Qualitative evaluations. In-domain: fastMRI. Out-of-domain: Stanford FSE.
-Table 5: Out-of-domain evaluation among supervised and untrained methods and comparisons with self-validation-based early stopping[48].
-Method In-domain Out-domain Runtime (mean±$\pm$±std) PSNR SSIM PSNR SSIM Train Inference Trained U-Net 31.16 0.776 29.16 0.724 $\sim$1.5 days $0.1\pm 0.003$ sec Untrained CS-$\ell_{1}$ [18] 28.23 0.633 22.46 0.407 ZS-SSL [48] 32.00 0.773 31.74 0.805 $26.1\pm 3.5$ mins DIP [45] 29.16 0.628 28.89 0.664 $9.2\pm 0.3$ mins $\mathbf{A_{2}\mathunderscore_{64}}$ 27.62 0.575 26.03 0.550 $5.5\pm 0.1$ mins $\mathbf{A_{2}\mathunderscore_{64}}$ w.
-Self-Val 29.59 0.695 27.59 0.641 $0.2\pm 0.2$ mins $\mathbf{A_{2}\mathunderscore_{64}}$ (Ours) 32.07 0.781 31.43 0.790 $6.4\pm 0.4$ mins $\mathbf{A_{2}\mathunderscore_{64}}$ (Ours) w.
-Self-Val.
-31.97 0.776 31.30 0.800 $0.3\pm 0.1$ mins
+Table 5: Out-of-domain evaluation among supervised and untrained methods and comparisons with a self-validation-based early stopping strategy [46].
+Method In-domain Out-domain Runtime (mean±$\pm$±std) PSNR SSIM PSNR SSIM Train Inference Trained U-Net 31.16 0.776 29.16 0.724 $\sim$1.5 days $0.1\pm 0.003$ sec Untrained CS-$\ell_{1}$ [16] 28.23 0.633 22.46 0.407 ZS-SSL [46] 32.00 0.773 31.74 0.805 $26.1\pm 3.5$ mins DIP [43] 29.16 0.628 28.89 0.664 $9.2\pm 0.3$ mins $\mathbf{A_{2}\mathunderscore_{64}}$ 27.62 0.575 26.03 0.550 $5.5\pm 0.1$ mins $\mathbf{A_{2}\mathunderscore_{64}}$ + Early Stopped 29.59 0.695 27.59 0.641 $0.2\pm 0.2$ mins $\mathbf{A_{2}\mathunderscore_{64}}$ (Ours) 32.07 0.781 31.43 0.790 $6.4\pm 0.4$ mins $\mathbf{A_{2}\mathunderscore_{64}}$ (Ours) + Early Stopped 31.97 0.776 31.30 0.800 $0.3\pm 0.1$ mins
 4.4 Benchmark Results
-We then compare our regularized network priors with several established MRI reconstruction methods, including a supervised baseline, a state-of-the-art self-supervised method (ZS-SSL) [48] and other untrained networks (ConvDecoder [9], Deep Decoder [14]).
-We adopt two lightweight architectures with distinct configurations as the base model.
-Visual comparisons are shown in Fig.
+We adopt two most lightweight architectures as the base models, and compare our regularized network priors with several established MRI reconstruction methods, including a supervised baseline, a state-of-the-art self-supervised method based on unrolling networks (ZS-SSL) [46] and underparameterized untrained networks (ConvDecoder [6], Deep Decoder [12]).
+Visual comparisons are presented in Fig.
 7.
-Our methods enable the previously underperforming architectures to match the performance of ZS-SSL and the supervised UNet on in-domain fastMRI knee data (Tab.
-4) and surpass the trained UNet on out-of-domain Stanford 3D FSE data (Tab.
-5) as untrained methods are agnostic to domain shifts.
-The datasets for training the UNet contain 367 knee slices and 651 brain slices, respectively, covering various kinds of contrast weightings.
-Our enhanced networks also clearly outperform other untrained networks and the compressed-sensing (CS) based reconstruction.
-These results demonstrate the benefits of enhancing the untrained methods for generalizable reconstruction.
-ZS-SSL is an unrolling-based method where the network, i.e., a ResNet [44], is adopted as a denoiser.
-Comparisons with its UNet-variant, i.e.,ZS-SSL-UNet, are included in the appendix.
-Additionally, ZS-SSL uses a self-validation strategy for early stopping.
-We further compare it with our methods.
-Tab.
-5 shows that self-validation is not as effective as ours.
-This is because, while self-validation mitigates overfitting by stopping near the peak PSNR, it does not fundamentally address the issues of underperforming architectures, i.e., the peak PSNR remains subpar.
-Our methods alleviate overfitting while generally enhancing the peak performance (e.g., Fig.
-3), which we further analyze in the next section.
-Hence, self-validation-based early stopping is complementary to our goal of reducing the architectural sensitivity, and can be integrated into our approach to further shorten the reconstruction time as shown in Tab.
-5.
-Overall, our enhanced network priors achieve comparable performance than ZS-SSL while being orders of magnitude faster.
+Comparisons with state-of-the-arts.
+Our method enables the previously underperforming architectures to match the performance of ZS-SSL and that of the supervised UNet on fastMRI knee data (Tab.
+4), and surpass the trained UNet on out-of-domain Stanford 3D FSE data (Tab.
+5), demonstrating their advantages in generalizable reconstruction.
+Our enhanced networks also clearly outperform other lightweight untrained networks, i.e., ConvDecoder and Deep Decoder, which are designed to prevent overfitting.
+ZS-SSL is an unrolling method where the network, i.e., a ResNet [42], is adopted as a denoiser.
+Our enhanced network priors achieve comparable performance than ZS-SSL while being orders of magnitude faster, thanks to the more efficient base models.
+Comparisons with early stopping (ES). ZS-SSL uses a self-validation strategy for early stopping whereas our method does not necessitate ES.
+Figure 8: 8×8$\times$8 × undersampling
+Our method not only alleviates overfitting but more importantly, enhances the inter/extrapolation capabilities of the underperforming architectures, leading to higher peak performance than the original scheme (Tab.
+5).
+This cannot be achieved even with the best ES strategy, as ES only halts the training near the peak PSNR for a given architecture, but it cannot fundamentally improve the underperforming architectures whose peak PSNR remains subpar (Fig.
+8).
+We show in Tab.
+5 that self-validation based ES can be integrated into our approach to further shorten the reconstruction time from 6 $\rightarrow$0.3 mins.
+More examples are in Suppl.
+Fig.
+8.
 4.5 Spectral Bias Analysis
-Figure 8: Measurement of the spectral bias.
-Our methods allow high frequencies to be fitted at a later iteration, alleviating overfitting and generally beneficial to inpainting-like tasks (i.e., MRI) with higher peak performance.
-PSNR shown for $\mathbf{A_{2}\mathunderscore_{256}}$.
-To examine the efficacy of the proposed methods on frequency control over the network, we measure the spectral bias using the metric— frequency-band correspondence (FBC) [41] that calculates the element-wise $|\mathcal{F}(x)|/|\mathcal{F}(y)|$ between the output $x$ and target image $y$, categorizes it into five frequency bands radially and computes the per-band averages.
+Figure 9: Measurement of spectral bias.
+Underperforming architectures (e.g., $A_{2}\mathunderscore_{256}$, $A_{8}\mathunderscore_{256}$) tend to learn high frequencies hastily, overfit more easily, and extrapolate poorly.
+Our method effectively mitigates these shortcomings.
+To examine how the proposed methods influence the frequency bias of the network, we measure the spectral bias using the metric— frequency-band correspondence (FBC) [39], which first calculates the element-wise $|\mathcal{F}(x)|/|\mathcal{F}(y)|$ between the output $x$ and target image $y$, categorizes it into five frequency bands radially and then computes the per-band averages.
 Higher values indicate higher correspondence.
-We trace the evolutions of FBC for $\mathbf{A_{8}\mathunderscore_{256}}$ and $\mathbf{A_{2}\mathunderscore_{256}}$ and the corresponding PSNR curves throughout the training iterations in three different tasks.
-As shown in Fig.
-8, the original architectures tend to readily fit all frequencies, especially $\mathbf{A_{2}\mathunderscore_{256}}$, corresponding to its worst performance among all compared architectures.
-Our methods delay the convergence of higher frequencies in both architectures across all three tasks as designed, which generally leads to higher peak performance for inpainting-like tasks.
-We hypothesize that a bias towards lower frequencies helps the model leverage spatial information for improved inter/extrapolation capabilities.
-Qualitative results in Fig.
-9 validate our hypothesis.
-Given that MRI reconstruction resembles an inpainting task for $k$-space measurements, we expect similar improvement using our regularized network priors, as proved in our experiments.
-Note that although a stronger spectral bias may not always be optimal, the hyperparameters of our methods enable flexible adaptations for the specific task (more experiments in the appendix).
-Figure 9: Inpainting and denoising ($\sigma=25$) experiments for demonstrating the improvement on underperforming architectures.
+We trace the evolutions of FBC for $\mathbf{A_{8}\mathunderscore_{256}}$ and $\mathbf{A_{2}\mathunderscore_{256}}$ and the corresponding PSNR curves throughout the training iterations in three tasks.
+Fig.
+9 shows that the original underperforming architectures tend to fit all frequencies more readily, including high frequencies.
+This is more evident in $\mathbf{A_{2}\mathunderscore_{256}}$, corresponding to its worst performance among all compared architectures.
+Our methods substantially delay the convergence of higher frequencies for all three tasks as designed, which decouples the learning of different frequencies.
+This leads to prolonged denoising effects and enhanced performance for inpainting tasks, including MRI reconstruction, as qualitatively shown in Fig.
+10.
+We speculate that a stronger bias towards lower frequencies helps the model leverage better the spatial information which improves its inter/extrapolation capability.
+Given that MRI reconstruction resembles an inpainting task for $k$-space measurements, we expect similar improvements in natural image inpainting shown in Fig.
+10 with our regularized network priors for $k$-space interpolation in MRI experiments.
+Figure 10: Experiments on natural image inpainting and denoising ($\sigma=25$). Our method improves the extrapolation and denoising capabilities of underperforming architectures.
 5 Conclusion
 We introduce efficient, architecture-agnostic methods for frequency control over the network priors, offering a novel solution to simultaneously address the key challenges present in untrained image reconstruction.
-Our approach requires only minimal modifications to the original DIP scheme while achieving significant gains in accruacy and efficiency as evidenced in MRI reconstruction and natural image restoration tasks, making it a stronger zero-shot baseline reconstructor with the potential for seamless integration with other advancements in self-supervised reconstruction.
-Appendix
-A Natural image experiments
-A.1 Inpainting
-To the best of our knowledge, our method is the first attempt that addresses challenges related to architecture, overfitting and runtime simultaneously.
-To compare with prior DIP methods, we 1) first employ the network configurations that work best in their respective settings [7, 41], i.e., the architectures used are different in different methods, as shown in Fig.
-10 and Fig.
-11, we then 2) use an underperforming architecture for all competing methods, as shown in Fig.
-12.
-Figure 10: Qualitative comparisons with previous DIP methods on inpainting.
-Figure 11: Qualitative comparisons with previous DIP methods.
-Figure 12: When unsuitable architectures are used, e.g., UNet w.
-deconvolutions, ResNet w/o.
-upsampling, SGLD [7] fails to perform restoration.
-This confirms the importance of architectural decisions in DIP, and that previous methods do not address the architecture-related challenges.
-Note that deconvolutions have been reported to be not suitable for DIPs [30, 15].
-Similarly for ResNet, which does not have any upsampling layers [30].
-A.2 Denoising
-We first show the results when each method operates in its original setting (Fig. 13), and then evaluate them when an unsuitable architecture is used (Fig. 14).
-Figure 13: Qualitative comparisons with previous DIP methods on denoising. Ours trades off the metrics for sharpness.
-Figure 14: Denoising results. The base network used in each competing method is replaced by an underperforming architecture, i.e., $\mathbf{A_{2}\mathunderscore_{256}}$.
-A.2.1 Transformer.
-Figure 15: Qualitative denoising results of a transformer [4]. Our method substantially alleviates the overfitting and enhances the peak PSNR.
-Besides CNN, we show here the result on Swin U-Net [4], which consists of only Swin Transformer blocks and skip connections, i.e., no upsampling is involved.
-As noted in a recent study [30], the unlearnt upsampling is the driving force behind the spectral bias of DIP, and such transformers are more difficult to perform denoising.
-This finding assumes the white noise as the input.
-Here, we show that constraining the bandwidth of the noise input enables long-lasting denoising even in such models.
-B Influences of hyperparameters
-As our methods contain several hyperparameters, we visualize their influences on the frequency control and hence the regularization effects over the output image in Fig. 16.
-$s$ and $\sigma$ are associated with the Gaussian blur kernel applied on the noise input.
-The larger the $s$ and/or $\sigma$ is, the more high frequencies are removed from the noise input (i.e., the smoother the output is).
-$M$ and $\beta$ are for adjusting the attenuation extents of the Kaiser-based upsamplers.
-The higher the $M$ and/or $\beta$ is, the larger the attenuation of the high-frequency replica (i.e., the smoother the output is).
-Figure 16: Visualizations of the frequency control with our methods in denoising experiments.
-C Comparisons with early-stopping
-Figure 17: Comparisons with self-validation-based early stopping.
-Although early stopping alleviates overfitting, it cannot fundamentally improve the underperforming architectures.
-Our methods can improve their peak performance while also mitigating overfitting.
-To complement Table 5 in the main text, here we visualize in Fig.
-17 that early stopping, even though prevents further performance decay, cannot fundamentally improve the underperforming architectures.
-In other words, early stopping could not cope with ill-designed architectures.
-D Architectural influences
-To better inform our method design, we 1) investigate the architectural influences in the context of MRI reconstruction, and also 2) validate our findings on image inpainting and denoising.
-Our results confirm that the reconstruction outcome is sensitive to basic architectural properties.
-D.1 Crucial Architectural Elements
-We first pinpointed the core architecture elements that have a critical impact on the performance.
-Experimental setup i.
-Since a decoder is the minimum requirement for reconstruction, we experimented with two types of 7-layered decoder-only architectures, i.e., ConvDecoder [9] and Deep Decoder [14].
-Experiments were performed on the 4× under-sampled multi-coil knee MRI from fastMRI database [25].
-Figure 18: Influences of architectural elements. Results averaged across three different widths.
-Upsampling (interpolation filter).
-Fig.
-18(a) suggests an interesting result: removing the unlearnt upsampling, e.g., bilinear, leads to either failure or unstable results (see gray curves).
-Unlike transposed convolution, the unlearnt upsampler is essentially a fixed low-pass interpolation filter that attenuates the introduced high-frequency replica and also the signal.
-Frequency response of bilinear interpolation filter decays more rapidly than that of nearest neighbor as the frequency increases (Fig.
-18 (b)), suggesting stronger attenuation and smoothing effects.
-Hence, bilinear upsampling typically biases the network towards generating smoother outputs, as prevalent in generative models [40].
-Transposed convolutions, however, are not guaranteed to be low-passed as they are learnable.
-Due to the spectral bias of network layers, they may be low-passed during early training to still enable reconstruction, but the results could be unstable (green curves).
-Convolutional layers.
-When the unlearnt upsampling operations are absent, ConvDecoder ($3\times 3$) still enables reconstruction while Deep Decoder ($1\times 1$) fails completely (Fig.
-18(a)).
-A similar phenomenon is also reported in image denoising [5, 30].
-This again be attributed to CNN’s inherent spectral bias and suggests that the size of the kernel also matters, further corroborated in Tab.
-7.
-Discussion.
-Results of this pilot experiment suggest that the spatial kernels with low-pass characteristics, either learnable or unlearnt, are crucial to the success of untrained network priors.
-In particular, bilinear upsampling with a fixed low-pass filter produces more stable and better results (blue curves).
-D.2 Depth, Width and Skip Connections
-Here, we demonstrate that insights gained about the unlearnt upsampling can aid in understanding the connection between architectural characteristics and the reconstruction task.
-Experimental setup ii.
-For this large-scale validation, we experimented with an isotropic encoder-decoder architecture used in the original DIP, i.e., equal width and kernel size for all layers throughout the network.
-Design choices are detailed in Tab.
-6.
-Experiments were performed on the publicly available 4×4$\times$4 × under-sampled multi-coil knee MRI from fastMRI database [24].
-Table 6: Test bed for studying the architectural influences of an encoder-decoder untrained networks.
-Archi.
-Type Depth (d) # of Skips (s) Width (w) Kernel Size (k) $\mathbf{A_{d}\mathunderscore_{s}\mathunderscore_{w}\mathunderscore_{k}}$2-L, 3-L, 4-L, 5-L, 8-L} {zero, half, full} {32, 64, 128, 256} $3\times 3,5\times 5$}
-Table 7: Influences of typical architectural design choices in knee reconstruction.
-Deeper and/or Narrower architectures tend to perform better; skip connections influence the deep architectures more; larger kernels perform slightly better.
-$\mathbf{A_{8}\mathunderscore_{full}\mathunderscore_{32}\mathunderscore_{3}}$ performs the best (in lime); $\mathbf{A_{2}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ performs the worst (in red).
-Width ($\downarrow$) Depth ($\uparrow$) Archi.
-PSNR SSIM Archi.
-PSNR SSIM Archi.
-PSNR SSIM Archi.
-PSNR SSIM $\mathbf{A_{2}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ 26.67 0.530 $\mathbf{A_{2}\mathunderscore_{full}\mathunderscore_{128}\mathunderscore_{3}}$ 27.12 0.543 $\mathbf{A_{2}\mathunderscore_{full}\mathunderscore_{64}\mathunderscore_{3}}$ 27.70 0.583 $\mathbf{A_{2}\mathunderscore_{full}\mathunderscore_{32}\mathunderscore_{3}}$ 28.47 0.641 $\mathbf{A_{3}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ 28.22 0.590 $\mathbf{A_{3}\mathunderscore_{full}\mathunderscore_{128}\mathunderscore_{3}}$ 28.59 0.605 $\mathbf{A_{3}\mathunderscore_{full}\mathunderscore_{64}\mathunderscore_{3}}$ 28.55 0.616 $\mathbf{A_{3}\mathunderscore_{full}\mathunderscore_{32}\mathunderscore_{3}}$ 29.25 0.660 $\mathbf{A_{4}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ 28.68 0.617 $\mathbf{A_{4}\mathunderscore_{full}\mathunderscore_{128}\mathunderscore_{3}}$ 28.95 0.622 $\mathbf{A_{4}\mathunderscore_{full}\mathunderscore_{64}\mathunderscore_{3}}$ 28.87 0.624 $\mathbf{A_{4}\mathunderscore_{full}\mathunderscore_{32}\mathunderscore_{3}}$ 29.70 0.671 $\mathbf{A_{5}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ 28.61 0.613 $\mathbf{A_{5}\mathunderscore_{full}\mathunderscore_{128}\mathunderscore_{3}}$ 28.87 0.615 $\mathbf{A_{5}\mathunderscore_{full}\mathunderscore_{64}\mathunderscore_{3}}$ 29.33 0.648 $\mathbf{A_{5}\mathunderscore_{full}\mathunderscore_{32}\mathunderscore_{3}}$ 29.81 0.680 $\mathbf{A_{8}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ 28.98 0.625 $\mathbf{A_{8}\mathunderscore_{full}\mathunderscore_{128}\mathunderscore_{3}}$ 29.33 0.637 $\mathbf{A_{8}\mathunderscore_{full}\mathunderscore_{64}\mathunderscore_{3}}$ 29.45 0.651 $\mathbf{A_{8}\mathunderscore_{full}\mathunderscore_{32}\mathunderscore_{3}}$ 30.04 0.695 Skip Connections (--) Kernel Size ($\uparrow$) $\mathbf{A_{2}\mathunderscore_{half}\mathunderscore_{256}\mathunderscore_{3}}$ 26.91 0.535 $\mathbf{A_{2}\mathunderscore_{zero}\mathunderscore_{256}\mathunderscore_{3}}$ 26.83 0.535 $\mathbf{A_{2}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ 26.67 0.530 $\mathbf{A_{2}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{5}}$ 26.98 0.550 $\mathbf{A_{4}\mathunderscore_{half}\mathunderscore_{256}\mathunderscore_{3}}$ 28.55 0.621 $\mathbf{A_{4}\mathunderscore_{zero}\mathunderscore_{256}\mathunderscore_{3}}$ 27.54 0.697 $\mathbf{A_{5}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ 28.61 0.613 $\mathbf{A_{5}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{5}}$ 28.82 0.624 $\mathbf{A_{8}\mathunderscore_{half}\mathunderscore_{256}\mathunderscore_{3}}$ 29.12 0.669 $\mathbf{A_{8}\mathunderscore_{zero}\mathunderscore_{256}\mathunderscore_{3}}$ 28.51 0.609 $\mathbf{A_{8}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{3}}$ 28.98 0.625 $\mathbf{A_{8}\mathunderscore_{full}\mathunderscore_{256}\mathunderscore_{5}}$ 29.12 0.634
-Figure 19: Generalizability of different architectures on the masked regions.
-What do deeper and narrower architectures produce?
-(Tab.
-7).
-Theoretically, as the number of layers (depth) or channels (width) increases, the ability of the network to learn arbitrarily high frequencies (details, noise) is typically increased [38].
-While this is true for width, we have found that the effect on depth turns out to be attenuated by unlearnt upsampling.
-As evidenced in Fig.
-19, deeper architectures typically generate smoother images, exhibiting a stronger preference for low-frequency information, whereas shallower counterparts, even though they have fewer parameters, are more susceptible to noise and overfitting (red arrows).
-This is more evident when comparing the same architectures with just different upsamplers, where the architectures with bilinear upsampling (stronger attenuation) are less prone to overfitting than the ones using nearest neighbor (NN) upsampling (cyan vs.
-blue).
-Hence, it is not merely the number of parameters but the architectural characteristics promoting low frequencies that seem to be the primary reason for the high performance.
-Note that all these results are only achievable when unlearnt upsampling is involved (gray dashed curves).
-Skip connections.
-Deep architectures with zero skip connection converge more slowly and may lead to over-smoothing as shown in Fig.
-19 (red curves).
-Skip connections greatly alleviate this issue and introduce more details (cyan curves), which we speculate could be due to the "reduced effective up-sampling rate".
-Yet, excessive skip connections make a deep architecture behave similarly as a shallower one, generating more noise (Fig.
-19 right).
-Overall, they exert a greater influence on deeper architectures ($\mathbf{A_{8}\mathunderscore_{zero}}<\mathbf{A_{8}\mathunderscore_{full}}<\mathbf{A_{8}\mathunderscore_{half}}$) compared to shallower ones ($\mathbf{A_{2}\mathunderscore_{full}}\approx\mathbf{A_{2}\mathunderscore_{zero}}$).
-D.3 Validations on image denoising and inpainting.
-We reaffirmed our observations above on image denoising and inpainting, as shown in Fig. 20, Fig. 22, Fig. 23 and Fig. 21.
-Figure 20: Denoising experiments.
-(Left) In non-convolutional networks, removing the upsampling hampers the denoising capability, which cannot be compensated by merely adjusting the network to be more under-parameterized.
-Transposed convolutions result in a more rapid decline in performance than bilinear upsampling.
-(Right) Convolutional layers alone exhibit certain denoising effects but necessitate early stopping.
-The showcased image is from the classic dataset Set9 [8].
-Figure 21: Deep architectures with zero skip connection converge more slowly, i.e., $\mathbf{A_{8}\mathunderscore_{zero}\mathunderscore_{256}}$
-We argue that the understanding of the upsampling and its interactions with other architectural elements can help explain why deeper networks with fewer skip connections converge more slowly, generate smoother outputs and are less prone to overfitting.
-Concretely, the upsampling operation inserted in-between the decoder layer slows down the generation of high frequencies required for transforming the lower-resolution feature maps into the higher-resolution target image, primarily due to its role as a fixed low-pass filter.
-As the network depth increases, the degree of smoothness increases (Fig.
-22).
-Skip connections notably accelerate the convergence (Fig.
-21) and ameliorate the over-smoothing issue, likely due to the reduced "effective" upsampling rate.
-All these observations are consistent with our MRI experiments in Sec.
-7.
-Figure 22: Denoising experiments. Deeper architectures with few or no skip connections tend to generate smoother outputs compared to the shallower ones.
-Figure 23: Inpainting experiments.
-Deeper architectures with few or no skip connections tend to generate smoother predictions for the masked regions than the shallower architectures.
-Skip connections make deep architectures perform similarly as the shallower ones.
-E Comparisons with ZS-SSL-UNet
-In the main paper, we have included the results of the ResNet version of ZS-SSL [48].
-Here, we constructed a UNet variant of it, dubbed ZS-SSL-UNet.
-As shown in Tab.
-8, the architecture type impacts not only DIP but also deep unrolling networks, and potentially a broader area, which worth future investigations.
-Table 8: Quantitative evaluations. Runtime: mean $\pm$ std mins per slice.
-[…]
+Our approach requires only minimal modifications to the original DIP scheme while achieving significant gains in accruacy and efficiency as evidenced in MRI reconstruction and natural image restoration tasks, making it a stronger zero-shot reconstructior with the potential for seamless integration with other advancements in self-supervised reconstruction.
+Acknowledgements
+This work was supported in part by the United States National Institutes of Health (NIH) through grants R01CA266702 and R01EB035160.